### v1

# 11footnotetext: Equal Contribution

Abstract
Weak-to-Strong Generalization (Burns et al., 2024) is the phenomenon whereby a strong student, say GPT-4, learns a task from a weak teacher, say GPT-2, and ends up significantly outperforming the teacher. We show that this phenomenon does not require a strong learner like GPT-4.
We consider student and teacher that are random feature models, described by two-layer networks with a random and fixed bottom layer and trained top layer. A ‘weak’ teacher, with a small number of units (i.e. random features), is trained on the population, and a ‘strong’ student, with a much larger number of units (i.e. random features), is trained only on labels generated by the weak teacher. We demonstrate, prove, and understand how the student can outperform the teacher, even though trained only on data labeled by the teacher. We also explain how such weak-to-strong generalization is enabled by early stopping. Importantly, we also show the quantitative limits of weak-to-strong generalization in this model.
1 Introduction
Today’s large foundation models exhibit extremely strong capabilities. For many tasks of interest, their performance is on par with or exceeds that of humans. In the foreseeable future, it is likely that we will have models that are much stronger than humans in capabilities, possibly through a delicate combination of supervised learning and reinforcement learning at scale. This motivates the study of superalignment: Can we humans, who will be much weaker in intelligence than these models at that point, still be able to supervise them effectively to align their behaviors with human values?
Burns et al. (2024) recently introduced the concept of weak-to-strong learning as a simple setup for studying the problem of superalignment. To enable iterative empirical progress on this problem even before superhuman models are available, they consider a surrogate setting where a weak teacher model instead of a human is trying to supervise a strong student model. More specifically, they fine-tuned a weak teacher model, such as GPT-2 (Radford et al., 2019), on task-specific data with true labels, simulating a human who has mastered a specific task to a certain extent. They then used this teacher model to annotate data for the same task, and fine-tuned a strong student model, such as GPT-4 (Achiam et al., 2023), on the annotated data. Perhaps surprisingly, they demonstrated the so-called weak-to-strong generalization phenomenon: the student model consistently outperforms the teacher model, even though the labels generated by the teacher model are imperfect.
A plausible explanation is that the student model is so strong that it can understand the essence of the task from imperfect labels. But this behavior should not be taken for granted. A crucial component of the setup is early stopping, since training the student model for too long can make it overfit to the teacher’s errors, resulting in no improvement or even worse performance. Indeed, Burns et al. (2024) found that stopping training at the right time can significantly improve weak-to-strong generalization.
Nevertheless, it remains unclear how general this phenomenon is beyond the specific experiment setup in Burns et al. (2024) and what other factors contribute to weak-to-strong generalization. For example, does it require an advanced model like GPT-4 with many emergent capabilities, or can it also arise in much simpler models? If the model sizes are further scaled up, will weak-to-strong generalization be enhanced or diminished? Without understanding its underlying principles, relying on weak-to-strong generalization for superalignment may be risky.
Figure 1: Illustration of our setup for weak-to-strong generalization. The teacher model of smaller size is trained on ground truth labels, while the student model of larger size is trained with early stopping on labels produced by the teacher model.
In this paper, we demonstrate that weak-to-strong generalization can happen even in models as simple as “random feature models”, i.e., two-layer networks with a random bottom layer and only the top layer trained. More specifically, we study the setup where the teacher and student models are two-layer networks with $M_{\mathrm{TE}}$ and $M_{\mathrm{ST}}$ (hidden, random) units, respectively. The student model is significantly stronger than the teacher, in the sense that it has many more units than the teacher, $M_{\mathrm{ST}}\gg M_{\mathrm{TE}}$. We train the top layer weights of both networks while fixing the bottom layer weights after drawing them from the same spherical distribution. As in Burns et al. (2024), the teacher model is trained with true labels, and the student model is trained only with labels generated by the teacher as described in Figure 1. We further assume that the teacher model is trained so well that it attains the minimum possible population loss. Thus, the only remaining source of error is the approximation error due to the finite number of units. See Section 2 for details of the setup.
Our contributions.
We analyze and explain how weak-to-strong generalization can arise in random feature models, if the student model is trained by gradient descent with early stopping. We rigorously prove that the student model can have a significantly smaller error than the teacher model, even as small as a power of the teacher’s error, with an exponent strictly greater than $1$. We establish the following results, where $\mathcal{L}_{\mathrm{TE}}$ and ${\mathcal{L}_{\mathrm{ST}}}$ denote the population losses of the teacher and student models, respectively:
  1. 1.
    For two-layer ReLU networks we show that the gap between the models can be arbitrarily large, in the sense that the ratio $\mathcal{L}_{\mathrm{TE}}/{\mathcal{L}_{\mathrm{ST}}}$ can be arbitrarily small (i.e. the student model is better than its teacher by an arbitrarily large factor), and thus the Performance Gap Recovered (PGR) introduced by Burns et al. (2024, see definition in Section 2 Equation 9) can be arbitrarily close to 1. Furthermore, relying on the Gaussian Universality Ansatz, we show an even stronger gap, in the sense that the student learns polynomially better, with ${\mathcal{L}_{\mathrm{ST}}}=\tilde{O}\left(\mathcal{L}_{\mathrm{TE}}^{1.49}\right)$ (Theorem 3.1).
  2. 2.
    For two-layer linear networks learning linear target functions over a specific but simple anisotropic Gaussian input distribution, we show that ${\mathcal{L}_{\mathrm{ST}}}=\tilde{O}\left(\mathcal{L}_{\mathrm{TE}}^{2}\right)$ (Theorem 3.2).
  3. 3.
    We also study the limit of weak-to-strong generalizatoin and show that for any random feature models, as long as the teacher model is optimally trained, the quadratic gap in the linear network case is the largest possible, i.e., ${\mathcal{L}_{\mathrm{ST}}}=\Omega(\mathcal{L}_{\mathrm{TE}}^{2})$ (Theorem 4.1). In particular, this implies that weak-to-strong generalization cannot arbitrarily “boost” a very weak teacher, and that in order to get student error approaching zero, the teacher error must also approach zero.
  4. 4.
    We explain how this arises as a result of early stopping. Gradient descent with early stopping implicity shrinks higher order eigencomponents of the target, and thus helps the student to eliminate spurious noise in those components introduced by the coarseness of the weak model, taking advantage of the target being aligned with the leading components (Section 5).
More generally, we provide a framework to understand weak-to-strong generalization and the optimal early stopping time for a broad class of random feature models and target functions (Section 6).
1.1 Related work
Empirical Work.
Following up on Burns et al. (2024)’s work, several variants of weak-to-strong generalization have been studied in Shin et al. (2025); Somerstep et al. (2024) in supervised fine-tuning settings. Yang et al. (2024); Bansal et al. (2024) show that small models can be used to generate reasoning data with Chain-of-Thought (CoT) for supervising large models. In particular, Bansal et al. (2024) point out that this can be more compute-optimal as compared to generating the data from large models. A number of other works have explored weak-to-strong generalization in better aligning today’s LLMs (Ji et al., 2024; Tao & Li, 2024).
Theoretical Understanding of Weak-to-Strong Generalization.
Many existing theoretical works explicitly control the strong student’s expressive power to avoid fitting the weak teacher’s mistakes. Lang et al. (2024) proposed a theoretical framework that establishes weak-to-strong generalization when the strong student is assumed to only learn functions that are robust to mislabeled data among many correctly labeled neighbors. Based on this framework, Shin et al. (2025) further formulated a weak-to-strong generalization mechanism for data with easy and hard patterns, where the weak teacher correctly label some data by leveraging easy patterns, and the strong student further learns to also leverage hard patterns from these correctly labeled data. Another work by Charikar et al. (2024) studied in a general regression setup with squared loss, where a key example is training only the last linear layer of the teacher and student networks on population. They showed that if the strong student’s capacity is not large enough to express the weak teacher’s function, then the student can outperform the teacher by an amount quantified by how much the student does not fit the teacher’s labels. Mulgund & Pabbaraju (2025) further generalized this result to a broader class of loss functions, including the cross-entropy loss. However, all these results do not establish weak-to-strong generalization in a strict sense, since the strong students in these works are explicitly constrained to not be able to fit the teacher’s function, making it unclear whether the student is indeed a stronger model or just has a different inductive bias than the teacher. In contrast, the student in our work is clearly a much stronger model than the teacher as its larger width allows it to have a much richer representation power, and we rely on early stopping to implicitly regularize the student model to avoid fitting the teacher’s mistakes.
Wu & Sahai (2025) demonstrated that weak-to-strong learning can happen in an overparametrized spiked covariance model, where the student’s features are specified to form a certain spectrum that better captures the prior than the teacher. Somerstep et al. (2024) present a negative result on weak-to-strong generalization from a transfer-learning perspective, showing that naively fine-tuning the strong model leads to poor performance. Ildiz et al. (2024) show that in a linear regression setting training with weak labels can outperform training with strong labels under the same data budget but cannot improve the scaling laws. In contrast, our work demonstrates several natural cases in which weak-to-strong generalization yields strictly super-linear improvements over the teacher model.
Asymptotic Notation
We use $\Theta_{c}$, $\tilde{O}_{c}$, and $\Omega_{c}$, where $c$ is one or more variables, to indicate a multiplicative factor that depends on $c$. E.g., for any two quantities $\textit{exp}_{1},\textit{exp}_{2}$, that could depend on various variables including $d$ and $k$, writting $\textit{exp}_{1}=\tilde{O}_{d,k}(\textit{exp}_{2})$ means that there exists some function $f(d,k)$ and exponent $r\geq 0$ s.t. it always holds that $\textit{exp}_{1}\leq f(d,k)\cdot\textit{exp}_{2}\cdot\log^{r}(\textit{exp}_{2})$.
2 Two Layer Networks, Random Features, and the Weak-to-Strong Setup
We consdier learning a predictor $f:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{X}=\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$, aiming for small population loss $\mathcal{L}(f):=\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}[(f({\bm{x}})-f^{*}({\bm{x}}))^{2}]$, where $\mathcal{D}$ is some input distribution and $f^{*}:\mathcal{X}\rightarrow\mathcal{Y}$ our target function. Learning is done using two-layer networks with $m$ units of the form
$f_{{\bm{w}},{\bm{u}}}({\bm{x}})=\sum_{i=1}^{m}w_{i}\sigma(\langle{\bm{u}}_{i},{\bm{x}}\rangle),$ (1)
where $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is some fixed activation function (e.g. ReLU or linear), the bottom layer weights ${\bm{u}}_{i}\in\mathbb{R}^{d}$ are randomly drawn
[…]
from some fixed distribution $\mathcal{U}$ over $\mathbb{R}^{d}$ (e.g. spherical Gaussian or similar) and then fixed during training, and only the top layer weights $w_{i}\in\mathbb{R}$ are trained. Such models, which amount to learning a linear predictor over the random features defined by the outputs of the hidden units, have been extensively studied as finite approximations of kernel methodsArriaga & Vempala (2006); Rahimi & Recht (2007), as models and limits of deep learning in certain regimes Mei et al. (2021) and recently also for deriving scaling laws and optimal tradeoffs Lin et al. (2024); Paquette et al. (2024).
Weak-to-Strong Setup.
We train two networks, a ‘weak’ teacher network $f_{\text{teacher}}$ with $M_{\mathrm{TE}}$ units, and a ‘strong’ student network $f_{\text{student}}$ with $M_{\mathrm{ST}}\gg M_{\mathrm{TE}}$ units. First, we train the teacher network $f_{\text{teacher}}:=f_{{\bm{w}}_{t},{\bm{u}}_{t}}$ on the true target labels $(x,f^{*}(x))$ using very large amount of data. We will model this by training the teacher on population. That is, to train the teacher we draw the bottom layer weights $\{{\bm{u}}_{t,i}\}_{i=1}^{M_{\mathrm{TE}}}$ at random as in (2), and for most results in this paper we assume that the top layer weights achieves the population loss minimization (Assumption 2.1):
Assumption 2.1.
The teacher model attains the minimum loss among all possible models with the given bottom layers:
${\bm{w}}_{t}=\arg\min_{{\bm{w}}\in\mathbb{R}^{M}_{\mathrm{TE}}}\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}\left[\left(f_{{\bm{w}},{\bm{u}}_{t}}({\bm{x}})-f^{*}({\bm{x}})\right)^{2}\right].$ (3)
We denote the teacher’s population loss (which is the minimal possible loss using its $M_{\mathrm{TE}}$ random units) by
$\mathcal{L}_{\mathrm{TE}}:=\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}\left[\left(f_{\text{teacher}}({\bm{x}})-f^{*}({\bm{x}})\right)^{2}\right].$ (4)
We then train the student on inputs $x\sim\mathcal{D}$ but only using labels $f_{\text{teacher}}(x)$ generated by the teacher. To train the student, we draw the bottom layer at random according to (2), initialize the top layer weights to zero, $\forall_{i}{\bm{w}}_{s,i}(0)=0$, and train these weights using SGD on i.i.d. samples ${\bm{x}}(n)\sim\mathcal{D}$ labeled with by the teacher:
$w_{s,i}(n+1)=w_{s,i}(n)-\eta\cdot\frac{\partial}{\partial w_{s,i}}\!\left(f_{{\bm{w}}_{s}(n),{\bm{u}}_{s}}({\bm{x}}(n))-f_{\text{teacher}}({\bm{x}}(n))\right)^{2}.\\ \end{aligned}$ (5)
Gradient Flow Limit.
We assume that the student has unrestricted access to the examples labeled by the teacher, and so can take arbitrarily small learning rate $\eta$ and an appropriately large number of steps $N$. Accordingly, we study the behavior with infinitesimally small learning rate $\eta$ where the student’s dynamics (5) converge to the gradient flow trajectory:
$\frac{\mathrm{d}w_{s,i}(t)}{\mathrm{d}t}=-\frac{1}{M_{\mathrm{ST}}}\frac{\partial}{\partial w_{s,i}}\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}\!\!\left[\left(f_{{\bm{w}}_{s}(t),{\bm{u}}_{s}}({\bm{x}})-f_{\text{teacher}}({\bm{x}})\right)^{2}\right]$ (6)
where we replaced the step number $n$ with training time $t=n\eta M_{\mathrm{ST}}$. We add an additional factor of $M_{\mathrm{ST}}$ here to ensure the gradient flow has a proper limit when $M_{\mathrm{ST}}\to\infty$, which will be useful in the analysis of the rest of the paper. In our analysis, we consider a student trained with the gradient flow dynamics (6) with stopping time $T$ that needs to be specified and denote the resulting student predictor
$f_{\text{student}}=f_{{\bm{w}}_{s}(T),{\bm{u}}_{s}}.$ (7)
Infinitely Large Student.
Furthermore, for simplicity we consider an infinitely large student $M_{\mathrm{ST}}\to\infty$ (see formal mathematical definitions in Section 6). All of our infinite-width results can be extended to the finite-width case for sufficiently large width. $M_{\mathrm{ST}}$
Generalization Gap.
We evaluate the student on the ground truth labels (which they do not have access to!), that is
${\mathcal{L}_{\mathrm{ST}}}:=\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}[f_{\text{student}}({\bm{x}})-f^{*}({\bm{x}}))^{2}].$ (8)
We ask how the student error ${\mathcal{L}_{\mathrm{ST}}}$ compares with the teacher error $\mathcal{L}_{\mathrm{TE}}$ and, in particular, whether and how it can be smaller even though it only sees examples labeled by the teacher.
To quantify the extent to which the student can best its teacher, Burns et al. (2024) suggested the Performance Gap Recovered (PGR), which is defined as [^1] This definition in terms of loss is equal to Burns et al.’s definition in terms of accuracy.:
$\mathrm{PGR}=\frac{\mathcal{L}_{\mathrm{TE}}-{\mathcal{L}_{\mathrm{ST}}}}{\mathcal{L}_{\mathrm{TE}}-{\mathcal{L}_{\mathrm{ST}}}^{C}}\geq\frac{\mathcal{L}_{\mathrm{TE}}-{\mathcal{L}_{\mathrm{ST}}}}{\mathcal{L}_{\mathrm{TE}}}$ (9)
where ${\mathcal{L}_{\mathrm{ST}}}^{C}\geq 0$ is the “ceiling performance”(Burns et al., 2024) of the student (the loss the student model could potentially have attained with direct access to the real labels). We do not carefully define or analyze this ceiling performance, since we provide lower bounds on the right-hand side in (9), and hence on the $\mathrm{PGR}$ regardless of the ceiling.
Specific Activation and Data Models.
We study two different variants of the random feature model (aka 2-layer network with random and fixed bottom layer) introduced above, which differ in the choice of activation function and input distribution:
Model 2.2 ($2$-layer ReLU Network).
The activation functions $\sigma(z)=\max(z,0)$ is a standard ReLU function, the bottom layer weights ${\bm{u}}$ are uniform over the sphere, i.e. $\mathcal{U}=\text{Unif}(\mathbb{S}^{d-1})$ in (2), and the inputs ${\bm{x}}$ are also uniformly distributed on the sphere, i.e. $\mathcal{D}=\text{Unif}(\mathbb{S}^{d-1})$. In this model, we will consider target function $f^{*}(x)$ which are even polynomials.
Model 2.3 (Linear Network).
The activation functions $\sigma(z)=z$ is linear (i.e. this is a linear neural net), and the bottom layer weights ${\bm{u}}$ are standard Gaussian, i.e. $\mathcal{U}=N(0,{\bm{I}}_{d})$ in (2). This time the inputs ${\bm{x}}$ are non-isotropic and are drawn from a Gaussian $\mathcal{D}=N(0,\Psi)$ with diagonal covariance $\Psi=\text{diag}(\psi_{1},\psi_{2},\dots,\psi_{d})$ with $\psi_{i}$ decreasing. Here, we will consider linear targets supported on the first top coordinates.
2.3 is fairly general and was studied by e.g. Lin et al. (2024) as a model for studying scaling laws. It can also be thought of as capturing a model where each unit outputs $\langle{\bm{u}},\phi({\bm{x}})\rangle$, where $\phi(\cdot)$ captures what happens in additional lower layers, which creates a non-isotropic representation at the top layer.
3 Quantifying Weak-to-Strong Generalization
We are now ready to establish weak-to-strong generalization, starting with the ReLU 2.2:
Theorem 3.1 (Weak-to-Strong generalization with $2$-layer ReLU Network).
Consider the ReLU 2.2 and the weak-to-strong setup of Section 2, for any dimension $d$ and teacher size $M_{\mathrm{TE}}$. Consider a target $f^{*}$ that is an even polynomial [^2] Or more generally, a sum of a linear function and an even polynomial. of degree at most $k$ normalized s.t. $\mathbb{E}[{f^{*}}^{2}]=1$, with any even degree $k$. If $M_{\mathrm{TE}}\geq\Theta(d^{2k})$ and the student is trained with some stopping time $T=\Theta_{d,k}(\log M_{\mathrm{TE}})$, then we have that with probability at least $1-\frac{3}{M_{\mathrm{TE}}}$
${\mathcal{L}_{\mathrm{ST}}}\leq O_{d,k}\left(\frac{\log^{2}M_{\mathrm{TE}}+\log^{2}\mathcal{L}_{\mathrm{TE}}}{\sqrt{M_{\mathrm{TE}}}}\,\mathcal{L}_{\mathrm{TE}}\right),$
and so ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}\to 0$ and $\mathrm{PGR}\to 1$ as $M_{\mathrm{TE}}$ increases.
Furthermore, under the Gaussian Universality Ansatz [^3] The Gaussian Universality Ansatz states that when sampling ${\bm{x}}$ the Gaussian universality holds for the eigenfunctions in the sense that the expected risk remains unchanged if we replace them with Gaussian with appropriate parameters Simon et al. (2021). In our setting, this can be differently stated as the expected risk for 2.2 will remain unchanged if we replace it with infinitely dimensional 2.3 with the appropriate covariance. Formally, the lower bound on $\mathcal{L}_{\mathrm{TE}}$ and an upper bound on $T$ are conditional on the eigenframework. This assumption has been empirically verified in a number of settings (Simon et al., 2021; Canatar et al., 2021; Wei et al., 2022; Misiakiewicz & Saeed, 2024)., for [^4] This choice of dimension is arbitrary. We can get close to ${\mathcal{L}_{\mathrm{ST}}}=\tilde{O}_{d,k}(\mathcal{L}_{\mathrm{TE}}^{1.5})$ with a larger $d$. $d>200$, $\mathcal{L}_{\mathrm{TE}}=\Omega_{d,k}(\frac{1}{M_{\mathrm{TE}}^{1.02}})$, and so as $M_{\mathrm{TE}}\to\infty$ we have that
${\mathcal{L}_{\mathrm{ST}}}=\tilde{O}_{d,k}\left(\mathcal{L}_{\mathrm{TE}}^{1.49}\right).$
For a more detailed statement spelling out the dependnce on the degree $k$ and stopping time $T$ see Theorem 6.7 and for the proof see Appendix E.
Theorem 3.1 already establishes that in 2.2, weak-to-strong generalization provably happens for a large set of targets, and for large teachers, we can get a performance gap recovered (PGR) arbitrarily close to $1$. We see also that the PGR is not a fine enough measure of weak-to-strong generalization, since the separation is obtained when the teacher error is small, and the PGR does not tell us how small the student error ${\mathcal{L}_{\mathrm{ST}}}$ can be as a function of the teacher error $\mathcal{L}_{\mathrm{TE}}$: having $PGR\rightarrow 1$ only says that the student error ${\mathcal{L}_{\mathrm{ST}}}$ goes to zero superlinearly faster than the teacher error $\mathcal{L}_{\mathrm{TE}}$, but not how much faster. Indeed, we can see that even in the ReLU Model 2.2, the student error is polynomially smaller, by an exponent of at least $1.49$. This last result relies on the widely verified Gaussian Universality Ansatz, and is also verified in our experiments: in Figure 2(a), we show that when the target $f^{*}$ is linear, with proper early stopping time, the loss ratio ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}$ decreases when $M_{\mathrm{TE}}$ grows. Furthermore, in Figure 2(b), we observe the student loss ${\mathcal{L}_{\mathrm{ST}}}$ is polynomially smaller than the teacher loss $\mathcal{L}_{\mathrm{TE}}$ with an estimated exponent even above our bound. Indeed, we expect our bound on the behavior of ReLU Random Feaure Networks is loose, and it might be possible to establish an even stronger separation.
(a) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}$ v.s. $t$
(b) ${\mathcal{L}_{\mathrm{ST}}}$ v.s. $\mathcal{L}_{\mathrm{TE}}$
Figure 2: Weak-to-strong generalization happens in ReLU radnom feature networks (2.2) with input dimension $d=32$, student size $M_{\mathrm{ST}}=16384$, and teacher size $M_{\mathrm{TE}}\in\{16,\ldots,256\}$. We consider a linear target function $f^{*}({\bm{x}})=\langle\beta,{\bm{x}}\rangle$ for unit norm some $\beta$. Figure 2(a) plots the ratio between student loss ${\mathcal{L}_{\mathrm{ST}}}$ and teacher loss $\mathcal{L}_{\mathrm{TE}}$, with varying teacher size $M_{\mathrm{TE}}$ and gradient flow training time $t$. With appropriate stopping time, we see a significant weak-to-strong generalization gain. This gain diminishes with overtraining and running gradient flow to convergence, the student mimics the teacher, has the same error, and does not excite weak-to-strong generalization. In Figure 2(b), we fit the minimal student loss ${\mathcal{L}_{\mathrm{ST}}}$ (at the optimal optimal stopping time for each teacher size) as a power law function of the student loss $\mathcal{L}_{\mathrm{TE}}$, confirming Theorem 3.1. See Appendix H for simulation details.
In order to demonstrate a stronger separation theoretically, and also avoid the Gaussian Universality Ansatz, we turn to the Linear Network 2.3:
Theorem 3.2 (Weak-to-Strong Generalization with Linear Network).
Consider the linear network 2.3, with $\Psi_{d}=(1,\dots,1,\psi_{*},\ldots,\psi_{*})$, where $1$ is repeated $k$ times and $\psi_{*}=(d-k)^{-2/3}$ is repeated $d-k$ times, and a linear target $f^{*}({\bm{x}})=\langle{\beta},{{\bm{x}}}\rangle$ supported by the first $k$ coordinates, i.e. $\beta_{i}=0$ for $i>k$ that is normalized $\mathbb{E}[f^{*2}]=1$. In the weak-to-strong setting of Section 2, for any $k$ and $d>k+Ck^{3}$(where $C$ is an absolute constant), and a teacher of size $M_{\mathrm{TE}}=(d-k)^{2/3}$, if the student is trained until time $T=\Theta(\log M_{\mathrm{TE}})$, we have that with probability at least $0.99$ for all $M_{\mathrm{TE}}$,
${\mathcal{L}_{\mathrm{ST}}}\leq\tilde{O}(k\mathcal{L}_{\mathrm{TE}}^{2}).$
In particular, for any fixed $k$, as $d,M_{\mathrm{TE}}\rightarrow\infty$, we have that $\mathrm{PGR}\to 1$.
A more general statement for 2.3 that holds for any covaraiance matrix $\Psi$ and target is given in Theorem 6.8. For the proof, see Appendix D. We indeed observe the behavior described by the Theorem in simulation experiments, as presented in Figure 3, where we can see the student loss ${\mathcal{L}_{\mathrm{ST}}}$ is indeed quadratically smaller than the teacher loss $\mathcal{L}_{\mathrm{TE}}$, already for moderately sized networks.
(a) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}^{2}$ v.s. $t$
(b) ${\mathcal{L}_{\mathrm{ST}}}$ v.s. $\mathcal{L}_{\mathrm{TE}}$
Figure 3: Weak-to-Strong generalization happens in antitropic random linear feature networks (2.3). Here we used an input distribution as in Theorem 3.2, with $k=1$ and a target function $f^{*}=\langle e_{1},{\bm{x}}\rangle$ where $e_{1}$ is the first standard basis vector. Figure 3(a) plots the ratio between the student loss ${\mathcal{L}_{\mathrm{ST}}}$ and squared teacher loss $\mathcal{L}_{\mathrm{TE}}^{2}$, with varying teacher size $M_{\mathrm{TE}}$, and where the dimensionality $d=M_{\mathrm{TE}}^{3/2}$ as set as in the scaling of Theorem 3.2, as a function of the gradient flow time $t$. With proper early stopping time ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}^{2}$ converges to approximately $1$ as $M_{\mathrm{TE}}$ grows, confirming that for large $M_{\mathrm{TE}}$ we have ${\mathcal{L}_{\mathrm{ST}}}\propto\mathcal{L}_{\mathrm{TE}}^{2}$ as in Theorem 3.2. This is also confirmed in Figure 3(b), where we fit the student loss ${\mathcal{L}_{\mathrm{ST}}}$ as a power law function of teacher loss $\mathcal{L}_{\mathrm{TE}}$, and recover an excellent fit with an exponent very close to $2$. We again see overtraining diminishes weak-to-strong generalization. See Appendix H for simulation details.
3.1 $\Theta(1)$-error asymptotics
Both in Theorem 3.1 and Theorem 3.2, we get decaying teacher error $\mathcal{L}_{\mathrm{TE}}$ and hence $\text{PGR}\to 1$ as $M_{\mathrm{ST}}$ increases. This is in contrast to how PGR behaves with language tasks - in Burns et al. (2024) PGR does not get close to $1$ but is fluctuating between $20\%$ and $80\%$. In this subsection, we show an example of a linear network 2.3 where the teacher loss $\mathcal{L}_{\mathrm{TE}}$ behaves as constant as the number of teacher features $M_{\mathrm{ST}}$ increases. Additionally, we show that in 2.2 we can’t have $\Theta(1)$-error asymptotics in fixed dimension $d$.
Theorem 3.3 ($\Theta(1)$-error asymptotics).
If in 2.3 we take $f^{*}=x_{1}$, $\Psi_{d}=(1,\sqrt{\frac{\alpha}{d-1}},\dots,\sqrt{\frac{\alpha}{d-1}})$, and $M_{\mathrm{TE}}=\sqrt{\frac{d-1}{\alpha}}$, where $\alpha\in(\Theta(\frac{1}{d^{0.49}}),1)$ is a fixed constant, then for all $d$ with probability at least $0.99$
$\mathcal{L}_{\mathrm{TE}}>0.98\frac{\alpha}{1+\alpha}~{}\text{ and }~{}{\mathcal{L}_{\mathrm{ST}}}\leq 100\mathcal{L}_{\mathrm{TE}}^{2}.$
Note that in this case PGR can get close to $1$ for the right choice of $\alpha$ even as $\mathcal{L}_{\mathrm{TE}}$ and ${\mathcal{L}_{\mathrm{ST}}}$ are both bounded away from 00, even when $M_{\mathrm{TE}}\to\infty$. For the proof of Theorem 3.3 see Section D.2.
On the other hand, for ReLU 2.2 in fixed dimension $d$, we don’t have $\Theta(1)$-error asymptotics for any target function that is a sum of an even and a linear functions (i.e. that can be represented by 2.2).
Theorem 3.4 (No $\Theta(1)$-error asymptotics for ReLU ).
Under Gaussian Universality, if $f^{*}$ is spanned by the spherical harmonics of order that is even and at mos $k$ or $1$, for 2.2 in fixed dimension $d$
$\mathcal{L}_{\mathrm{TE}}=o(1),$
i.e. we have to have $\mathcal{L}_{\mathrm{TE}}\to 0$ as $M_{\mathrm{TE}}\to\infty$.
For the proof of Theorem 3.4 see Section G.2.
4 The Limit of Weak to Strong Generalization
One might ask if in some cases the student can achieve an even larger improvement upon the teacher than the quadratic bounds we have shown in Theorem 3.2. However, in Theorem 4.1 we show that for fairly general weak-to-strong setups, quadratic improvements are still the best that we could hope for. More specifically, for any two-layer model as in (1) with fixed random bottom layer as in (2), with any activation function and any bottom layer distribution $\mathcal{U}$, and any number of teacher units $M_{\mathrm{TE}}$, and as long as the teacher is optimally trained as in (3), for any student stopping time $T$, the student error can be at most quadaratically better than the teacher’s. That is, Theorem 4.1 shows that the separation in Theorem 3.2 is the largest possible. This result can hold even if the features are not i.i.d. sampled, see Theorem 6.11 for a result that applies in this case.
Theorem 4.1 (Limit of Random Feature Weak to Strong Generalization).
For any Random Feature Network Weak-to-Strong training as in Section 2, any normalized target $\mathbb{E}[{f^{*}}^{2}]=1$, and any student stopping time $T$, we have that
${\mathcal{L}_{\mathrm{ST}}}\geq\frac{\left(\sqrt{1+3\mathcal{L}_{\mathrm{TE}}}-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}}{4}\geq\frac{3}{4}\mathcal{L}_{\mathrm{TE}}^{2}.$
See Theorem 6.11 for a more general version and Appendix F for the proof.
Theorem 4.1 informs us of the limit of weak-to-strong generalization in random feature networks trained with early-stopped gradient descent. Together with Theorems 3.1 and 3.2, we see that even a Random Feature Network can exhibit weak to strong generalization, with PGR arbitrarily close to one, but that we cannot have arbitrarily high teacher error and arbitrarily low student error. For example, we cannot have a situation in which the teacher error is very close to the null risk $\mathbb{E}[{f^{*}}^{2}]$, i.e. the teacher is barely learning, while the student error is low (if $\mathcal{L}_{\mathrm{TE}}\approx 1$ then ${\mathcal{L}_{\mathrm{ST}}}\approx 1$). That is, weak-to-strong learning cannot significantly “boost” a teacher that is only slightly better than chance. And, in order for the student error to go to zero, the teacher error must also go to zero. Additionally, this theorem shows that if the teacer is going to zero, the student error can go to zero at most quadratically faster. An important caveat here is that these limitations on weak-to-strong learning apply only for Random Feature Networks. Understanding how these limitations play out in Random Feature Networks prompts one to ask whether such limits exist also in more complex model or more generally.
5 How Does Weak-to-Strong Generalization Happen?
Linear Network.
To understand how the student can correct their weak teacher and obtain a lower error, let us first consider linear networks as in 2.3. In this case, both weak and strong models, $f_{{\bm{w}},{\bm{u}}}({\bm{x}})$ are linear functions of ${\bm{x}}$, and thus can be written as $f_{{\bm{w}},{\bm{u}}}({\bm{x}})=\langle{\beta},{{\bm{x}}}\rangle$ where $\beta={\bm{w}}{\bm{U}}$. First, let us understand the student’s learning rule. Learning a predictor $f$ with an early stopped gradient descent shrinks $f-f_{{\bm{w}},{\bm{u}}}$ in the direction of $\sigma(\langle{\bm{u}}_{i},x\rangle)=({\bm{U}}_{S}x)_{i}$ proportional to its variance. With $M_{\mathrm{ST}}\to\infty$, the shrinkage of $f-f_{{\bm{w}},{\bm{u}}}$ will be in directions $x_{i}$ proportional to theire variance $\psi_{i}$. This is because as $M_{\mathrm{ST}}\to\infty$ we have that ${\bm{U}}_{S}^{\top}{\bm{U}}_{S}\to M_{\mathrm{ST}}I$. This makes learning directions $x_{i}$ with larger $\psi_{i}$, i.e. for smaller $i$, significantly faster than the ones with smaller $\psi_{i}$, i.e. for larger $i$. By appropriately choosing the stopping time, this allows us to effectively zero out the signal $f$ in directions $x_{i}$. Let us now turn to the teacher. Although the signal is entirely in the first few high energy coordinates of ${\bm{x}}$, if we have a small number of teacher nodes $M_{\mathrm{TE}}$, they will not be directly represented in the hidden layer, i.e. in ${\bm{U}}_{T}{\bm{x}}$. Instead, the teacher can only learn $\beta_{T}\in\text{span}({\bm{U}}_{T})$, and so learns $\beta_{T}$ to be the projection of $f^{*}$ to $\text{span}({\bm{U}}_{T})$. This means that $\beta_{T}$ has some energy (non-zero coefficients) in all the coordinates. The student shrinkage will reduce the noise in low energy coordinates by zeroing out the coefficients along those directions. Therefore, the weak-to-strong improvement comes exactly from improvement along the noise directions. Note that if we train for too long, the effect will be the same on all directions, so there will be no weak-to-strong generalization. So early stopping is crucial in this setup.
As student training progresses we fit more of the teacher’s signal, but the shrinkage effect reduces. This can be seen in Figures 2 and 3, where overtraining hurts weak-to-strong generalization. At the limit, as training time goes to infinity, the student would converge to exactly mimicking the teacher, and thus converge to the teacher error, without any weak-to-strong benefit, as can be clearly seen in Figure 2.
The difficulty of applying this reasoning is that we need to establish that a significant portion of the teacher error is actually present in the noise direction. In Theorem 3.1 and Theorem 3.2, we not only show that a significant portion of the teacher’s error $\mathcal{L}_{\mathrm{TE}}$ will be along the low energy directions but also that when the number of teacher’s units $M_{\mathrm{TE}}$ is large, most of the teacher’s error lies along the low energy, i.e. noise directions. Our main result, Theorem 6.3, handles this difficulty. Proving that most of the teacher error $\mathcal{L}_{\mathrm{TE}}$ is in the low energy directions allows us to conclude that the PGR goes to $1$ as the number of teacher units $M_{\mathrm{TE}}$ increases, instead of just being positive, and establish a quantitative relationship between the student and teacher errors.
$2$-layer ReLU
For ReLU, we have a similar situation, except now the relevant basis is not the coordinates of the input ${\bm{x}}_{i}$ but rather spherical harmonics. SGD shrinkage will be along a specific basis of functions with respect to the probability distribution $\mathcal{D}=\text{Unif}(\mathbb{S}^{d-1})$, which are the spherical harmonics $\{\{\phi_{i,k}\}_{i=1}^{N_{k}}\}_{k=1}^{\infty}$. The $i$ index indicates there are multiple spherical harmonics of order $k$, namely $N_{k}$ of them. So the teacher predictor is given by
$f_{\text{teacher}}({\bm{x}})=\sum_{i,k}\beta_{i,k}\phi_{i,k}({\bm{x}}).$
The student will learn the teacher’s coefficients for $\phi_{i,k}$, $\beta_{i,k}$, faster for the spherical harmonics of lower order $k$. So, if we select the stopping time $T$ correctly, the student will learn $\beta_{i,k}$ almost completely for all $k$ up to some index $k_{0}$ depending on $T$, while for $k>k_{0}$ the student will be very close to 00. This means that the student can denoise the direction of $\phi_{i,k}$ for $k>k_{0}$. So, in particular, if all of the signal $f^{*}$ is aligned with the directions of the first $k_{0}$ order of spherical harmonics, then everything the teacher learns in the directions $\phi_{i,k}$ for $k>k_{0}$ is the incorrectly learned noise. Zeroing out this incorrectly learned noise enables weak-to-strong generalization.
To understand why this is the case, we turn to the dynamics of GF (Equation 6) in Section 6. This will be simplified if we consider the student’s network as a kernel.
6 General Setup and Results
In this section, we state our result in its full generality. We first introduce our general setup and assumptions, and then state our main result on weak-to-strong generalization. Then we provide a negative result on the limit of weak-to-strong generalization.
6.1 Setup
Gradient Flow Dynamics.
We first introduce the function space view of gradient flow over the top layer of the student’s network Equation 6. The main benefit of working in function space from parameter space is allowing us to analyze the infinite-width limit in the same space. We first recall we parametrize the student in Equation 7 as:
$f_{\text{student}}=f_{{\bm{w}}_{s},{\bm{u}}_{s}}({\bm{x}})=\sum_{i=1}^{M_{\mathrm{ST}}}w_{s,i}\sigma(\langle{\bm{u}}_{s,i},{\bm{x}}\rangle).$
Let $\langle{\cdot},{\cdot}\rangle_{\mathcal{D}}$ be the inner product with respect to $\mathcal{D}$, defined as as $\langle f,g\rangle_{\mathcal{D}}:=\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}\left[f({\bm{x}})g({\bm{x}})\right]$. Let $\sigma_{s,i}({\bm{x}})=\sigma(\langle{\bm{u}}_{s,i},{\bm{x}}\rangle)$. Then, Equation 6 in function space is
$\frac{\mathrm{d}}{\mathrm{d}t}f_{{\bm{w}}_{s}(t),{\bm{u}}_{s}}({\bm{x}})=&-\frac{1}{M_{\mathrm{ST}}}\sum_{i=1}^{M_{\mathrm{ST}}}\mathbb{E}_{{\bm{y}}\sim\mathcal{D}}\langle{f_{{\bm{w}}_{s}(t),{\bm{u}}_{s}}({\bm{y}})-f_{\text{teacher}}({\bm{y}})},{\sigma_{s,i}({\bm{y}})}\rangle~{}\sigma_{s,i}({\bm{x}})\\ =&-\mathbb{E}_{{\bm{y}}\sim\mathcal{D}}\left[K_{M_{\mathrm{ST}}}({\bm{x}},{\bm{y}})(f_{{\bm{w}}_{s}(t),{\bm{u}}_{s}}({\bm{y}})-f_{\text{teacher}}({\bm{y}}))\right]\end{aligned}$ (10)
where $K_{M_{\mathrm{ST}}}({\bm{x}},{\bm{y}})=\frac{1}{M_{\mathrm{ST}}}\sum_{i=1}^{M_{\mathrm{ST}}}\sigma_{s,i}({\bm{x}})\sigma_{s,i}({\bm{y}})$ is the empirical kernel.
Taking $M_{\mathrm{ST}}\to\infty$, the solution of Equation 10 converges to the solution of the following equation:
$\frac{\mathrm{d}}{\mathrm{d}t}f_{t}({\bm{x}})=-\mathbb{E}_{{\bm{y}}\sim\mathcal{D}}\left[{\mathcal{K}}({\bm{x}},{\bm{y}})\left(f_{t}({\bm{y}})-f_{\text{teacher}}({\bm{y}})\right)\right],\end{aligned}$ (11)
where
${\mathcal{K}}({\bm{x}},{\bm{x}}^{\prime}):=\mathbb{E}_{{\bm{u}}\sim\text{unif}(\mathbb{S}^{d-1})}\left[\sigma(\langle{\bm{u}},{\bm{x}}\rangle)\sigma(\langle{\bm{u}},{\bm{x}}^{\prime}\rangle)\right].\end{aligned}$ (12)
The kernel ${\mathcal{K}}$ can be decomposed as ${\mathcal{K}}({\bm{x}},{\bm{x}}^{\prime})=\sum_{k\geq 1}\lambda_{k}e_{k}({\bm{x}})e_{k}({\bm{x}}^{\prime})$, where $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\cdots$ are the eigenvalues of the associated kernel operator in descending order, and $e_{1}({\bm{x}}),e_{2}({\bm{x}}),e_{3}({\bm{x}}),\dots$ are orthonormal eigenfunctions, that is, $\langle{e_{i}},{e_{j}}\rangle_{\mathcal{D}}=\delta_{ij}$. This allows us to decompose the gradient flow dynamics (Equation 10) into simple ODEs for each eigendirection, which can be easily solved. With this decomposition, we can derive the following closed-form expression for the gradient flow dynamics of the student (Equation 6) at time $t$:
$f_{t}=\sum_{k\geq 1}\left(1-e^{-\lambda_{k}t}\right)\langle{f_{\text{teacher}}},{e_{k}}\rangle_{\mathcal{D}}e_{k}$ (13)
We note that given data distribution $\mathcal{D}$ (and thus the inner-product $\langle{\cdot},{\cdot}\rangle_{\mathcal{D}}$), the above solution of gradient flow at time $t$ is only a function of target $f^{*}$, time $t$, and the kernel $\mathcal{K}$, which we denote by
$f_{t}=\mathcal{T}^{\mathcal{K}}_{t}(f_{\text{teacher}}),\quad\text{where}\quad\mathcal{T}^{\mathcal{K}}_{t}\triangleq\mathtt{id}-e^{-t\mathcal{K}},$ (14)
where $\mathtt{id}$ is the identity mapping. In other words, if different two-layer networks have different bottom layers but with the same induced kernel, they lead to same training behavior under gradient flow. For an outline of the derivation of gradient flow dynamics, see Appendix A.
Using Equation 13, we can understand how weak-to-strong happens from Section 5 more clearly. To instantiate Equation 13 to Section 5, note that $\langle f_{\text{teacher}},e_{k}\rangle_{\mathcal{D}}$ is what we call $\beta_{i}$ and $e_{i}({\bm{x}})$ are $x_{i}$ in 2.3 and spherical harmonics $\phi_{i,k}({\bm{x}})$ in 2.2. Note that since $\lambda_{k}$ are decreasing, $1-e^{-\lambda_{k}t}$ is closer to $1$ for larger eigenvalues $\lambda_{i}$, i.e. smaller indices $i$. Therefore, if there is $K$ for which $\lambda_{K+1}$ is significantly smaller than $\lambda_{K}$, then for stopping time $T\propto(\frac{1}{\lambda_{K}},\frac{1}{\lambda_{K+1}})$, $e^{-\lambda_{i}T}$ are all going to be close to 00 for $i\leq K$, but $e^{-\lambda_{i}T}$ for $i\geq K+1$ will be close to $1$, effectively zeroing out $f_{\text{teacher}}$ in directions $e_{\geq K+1}$. This is the intuition behind Theorem 6.3, which we state now.
Notations.
For $S\geq K$, let ${\bm{A}}_{S},{\bm{B}}_{S}\in\mathbb{R}^{m\times m}$ be the Gram matrices of random units projected onto or out of the top-$S$ eigenspace of ${\mathcal{K}}$, i.e., $A_{S,i,j}:=\sum_{k=1}^{S}\langle{g_{i}},{e_{k}}\rangle_{\mathcal{D}}\langle{g_{j}},{e_{k}}\rangle_{\mathcal{D}}$, $B_{S,i,j}:=\sum_{k\geq S+1}\langle{g_{i}},{e_{k}}\rangle_{\mathcal{D}}\langle{g_{j}},{e_{k}}\rangle_{\mathcal{D}}$. Let $\kappa_{S}$ be the following key quantity for characterizing the weak-to-strong generalization gap, which will be useful in Theorem 6.3,
$\kappa_{S}:=\frac{1}{1+\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)}.$ (15)
$\kappa_{S}$ measures the alignment of the teacher’s features with the top-$S$ eigenspace of the student kernel. As explained in Section 5, this misalignment is crucial for enabling weak-to-strong generalization in this setup. We omit the subscript $S$ and write ${\bm{A}},{\bm{B}}$ and $\kappa$ when $S$ is clear from the context.
Furthermore, we assume that the student’s kernel ${\mathcal{K}}$ correctly captures the prior of the target function $f^{*}$ in the sense that $f^{*}$ is supported on the top-$K$ eigenspace of ${\mathcal{K}}$ with nonzero eigenvalues:
Assumption 6.1 (Ground Truth).
$f^{*}$ is supported on the top-$K$ eigenspace of ${\mathcal{K}}$ corresponding to nonzero eigenvalues, i.e., if $\lambda_{k}=0$ for $k\leq K$ then $\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}=0$ and $\forall k>K:\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}=0$.
Finally, we will assume that no two units in the first layer of the teacher will be linearly dependent among the first $K$ directions $e_{i}$.
Assumption 6.2.
$\mathrm{rank}({\bm{A}}_{K})=K$.
This assumption is satisfied almost surely for 2.2 and 2.3.
6.2 Provable Weak-to-Strong Generalization
6.2.1 Main Theorems
Theorem 6.3 (Weak-to-Strong Multiplicative Error Improvement).
For any groundtruth $f^{*}$, p.s.d. kernel $\mathcal{K}$ and positive integer $K$ satisfying Assumption 6.1 and any teacher model $f_{\text{teacher}}$ satisfying Assumption 6.2 and Assumption 2.1 ($f_{\text{teacher}}$ attaining minimum loss), then for any time $T>0$, then the test loss of student $f_{\text{student}}\triangleq\mathcal{T}^{\mathcal{K}}_{T}(f_{\text{teacher}})$ has the following upper bound:
${\mathcal{L}_{\mathrm{ST}}}$ $\leq\inf_{S\geq K}\bigg{\{}\left(1-(1-\lambda_{S+1}^{2}T^{2})\kappa_{S}\right)\mathcal{L}_{\mathrm{TE}}+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\|f^{*}\|_{\mathcal{D}}^{2}\bigg{\}}.$
Theorem 6.3 shows that if there is a large enough eigengap between $\lambda_{i}$ so that $\kappa_{S}$ is close to $1$, then we can select the early stopping time $T$ so that the student is only left with $(1-\kappa_{S})$ fraction of teacher’s loss. See Appendix C for the proof.
In Theorem 6.5, we instantiate Theorem 6.3 for a $2$-layer network with arbitrary activation, defined in 6.4.
Model 6.4 ($2$-layer Network).
Consider a $2$-layer network with an activation function $\sigma$ that is bounded on $[-1,1]$. Let the distribution $\mathcal{D}$ be ${\bm{x}}\sim\text{Unif}(\mathbb{S}^{d-1})$. The first layer parameters ${\bm{u}}_{i}$ are initialized randomly isotropically, i.e. ${\bm{u}}_{i}\sim\text{Unif}(\mathbb{S}^{d-1})$. Let $\{\sigma_{i}\}_{i=1}^{\infty}$ be the nonzero coefficients of the activation function $\sigma(\langle{\bm{x}},{\bm{t}}\rangle)$ in the basis of spherical harmonics, i.e. $\sigma(\langle{\bm{x}},{\bm{t}}\rangle)=\sum_{i,k}\sigma_{i}\phi_{i,k}({\bm{x}})\phi_{i,k}({\bm{t}})$, where the sum goes over only the nonzero $\sigma_{i}$. Here $\phi_{i,k}$ is the $i$-th spherical harmonic of order $k$ and is also the eigenbasis of the induced kernel (Equation 12) [^5] The existence of such a decomposition follows from the Funk-Hecke formula.
Theorem 6.5 (Weak-to-Strong Generalization with 2-layer NN).
Under 6.4, let Assumption 6.1 be satisfied for groundturth $f^{*}$, population kernel $\mathcal{K}$ (Equation 12) and positive integer $K$. If $M_{\mathrm{TE}}\geq K$ and the teacher attains the minimum loss (Assumption 2.1), then for any $\delta_{T}\in(0,1)$ and the student trained to stopping time $T=\frac{1}{\sigma_{K}^{2}}\log\frac{1}{\delta_{T}}$, with probability at least $1-\frac{2}{M_{\mathrm{TE}}}-K\exp(-\frac{c}{2}\frac{M_{\mathrm{TE}}}{K})$ over the randomness of teacher features, it holds that
${\mathcal{L}_{\mathrm{ST}}}\leq\inf_{S\geq K}\Biggl{\{}\frac{4(\sum_{k=1}^{S}\sigma_{k}^{2})(\sum_{k=S+1}^{\infty}\sigma_{k}^{2})\log^{2}M_{\mathrm{TE}}}{\sigma_{S}^{4}M_{\mathrm{TE}}}\,\mathcal{L}_{\mathrm{TE}}+\frac{\sigma_{S+1}^{4}}{\sigma_{K}^{4}}\left(\log\frac{1}{\delta_{T}}\right)^{2}\,\mathcal{L}_{\mathrm{TE}}\Biggr{\}}+\frac{\delta_{T}}{2-\delta_{T}}\,\bigl{\|}f^{*}\bigr{\|}_{\mathcal{D}}^{2}.$
For the proof of this theorem see Appendix E.
Having established Theorem 6.5 for general 2-layer networks, we are ready to apply it to ReLU and linear networks, i.e., 2.2 and 2.3.
6.2.2 ReLU Networks
For ReLU activation in 2.2, the expansion of $\sigma$ in the basis of spherical harmonics $\phi_{k,i}$ will be $(\sigma_{1},\dots,\sigma_{1},\sigma_{2},\dots,\sigma_{2},\dots)$ where each $\sigma_{i}$ repeats $N_{k}=\frac{(2k+d-2)(k+d-3)!}{k!(d-2)!}$ times, which is the order of the $k$-th spherical harmonics in $d$ dimensions. Furthermore, $\sigma_{k}$ is nonzero only for even $k$ and $k=1$.
As $2$-layer ReLU networks, 2.2, can only represent functions spanned by spherical harmonics that are either order $1$ or even order, i.e. sums of linear and even functions, we make the following restrictions on the target function.
Assumption 6.6 (Ground Truth For $2$-layer ReLU).
$f^{*}$ is supported on the spherical harmonics of order $1$ or even order less than or equal to $k$. In this case, we will denote $K=N_{0}+N_{1}+\dots+N_{k}$.
Theorem 6.7 (Weak-to-Strong Generalization with 2-layer ReLU NN).
Under 2.2, let Assumption 6.6 be satisfied for $f^{*}$ and $k$ and $K$. If the teacher attains minimum loss $\mathcal{L}_{\mathrm{TE}}=\hat{\mathcal{L}}_{\min}$, then the student trained to early stopping time $T=\frac{1}{\sigma_{k}^{2}}\log\frac{1}{\delta_{T}}$ will have with probability at least $1-\frac{2}{M_{\mathrm{TE}}}-K\exp(-c\frac{M_{\mathrm{TE}}}{K})$,
${\mathcal{L}_{\mathrm{ST}}}\leq O_{d}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\cdot\mathcal{L}_{\mathrm{TE}}+O_{d}\left(\frac{1}{\sigma_{k}^{4}}\frac{1}{\sqrt{M_{\mathrm{TE}}}}\right)\cdot\left(\log\frac{1}{\delta_{T}}\right)^{2}\mathcal{L}_{\mathrm{TE}}+\frac{\delta_{T}}{2-\delta_{T}}\|f^{*}\|_{\mathcal{D}}^{2}.$
Here $c$ is an absolute constant.
For the proof of this theorem see Section E.1.
6.2.3 Linear Networks
Finally, we can state the general bound for linear networks, 2.3.
Theorem 6.8 (Weak-to-Strong Generalization in Linear Networks).
For 2.3, if $f^{*}$ is spanned by the first $K$ coordinates and if the teacher attains minimum loss $\mathcal{L}_{\mathrm{TE}}=\hat{\mathcal{L}}_{\min}$, then for the student trained until time $T$ we have with probability at least $1-\frac{4}{m}$ that
${\mathcal{L}_{\mathrm{ST}}}$ $\leq\inf_{S\geq K}\Biggl{\{}\frac{\bigl{(}\sum_{i=1}^{S}\psi_{i}\bigr{)}\bigl{(}\sum_{i=S+1}^{d}\psi_{i}\bigr{)}}{\psi_{S}^{2}}\,\frac{(\log m)^{2}}{m}\,\mathcal{L}_{\mathrm{TE}}+\psi_{S+1}\,T^{2}\,\mathcal{L}_{\mathrm{TE}}\Biggr{\}}+\frac{e^{-\psi_{K}T}}{2-e^{-\psi_{K}T}}\,\bigl{\|}f^{*}\bigr{\|}_{\mathcal{D}}^{2}.$
For the proof, see Appendix D.
6.3 Limits of Weak-to-Strong Generalization
In this subsection, we will prove a more general lower bound for student loss in weak-to-strong setups, including gradient flow as a special case. We will first introduce the key concept, shrinking optimality (Definition 6.9) and show that the gradient flow solution is shrinkage optimal with respect to its target function (Lemma 6.10). Then we will state the main result under shrinking optimality, Theorem 6.11, which shows that the student can be at most quadratically better than the teacher in weak-to-strong generalization.
Definition 6.9 (Shrinking Optimality).
Given two functions $f,f^{\prime}:\mathbb{R}^{d}\to\mathbb{R}$, we say $f$ is shrinnking-optimal with respect to $f^{\prime}$ if for any $0\leq\alpha\leq 1$, $\mathbb{E}_{{\bm{x}}}((f({\bm{x}})-f^{\prime}({\bm{x}}))^{2})\leq\mathbb{E}_{{\bm{x}}}((\alpha f({\bm{x}})-f^{\prime}({\bm{x}}))^{2})$.
We will show that for gradient flow solutions, the learned predictor is shrinkage optimal with respect to the target.
Lemma 6.10 (Shrinkage Optimality of Gradient Flow Solutions).
For any positive semi-definite kernel $\mathcal{K}$, time $T>0$, and ground truth $f^{*}$, the gradient flow solution $f_{T}=\mathcal{T}^{\mathcal{K}}_{T}(f^{*})$ is shrinkage optimal with respect to $f^{*}$ in the sense of Definition 6.9.
Lemma 6.10 shows that in the setup considered in Section 2, the teacher and the student both satisfy Definition 6.9 with respect to target and the teacher predictor.
Theorem 6.11 (Genereal Limitation of Weak-to-Strong Generalization).
For any target $f^{*}$ with unit norm, i.e., $\|f^{*}\|_{\mathcal{D}}=1$, for any teacher $f_{\text{teacher}}$ that is shrinking-optimal to $f^{*}$ and any student $f_{\text{student}}$ that is shrinking-optimal to $f_{\text{teacher}}$, it holds that
${\mathcal{L}_{\mathrm{ST}}}\geq\frac{\left(\sqrt{1+3\mathcal{L}_{\mathrm{TE}}}-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}}{4}\geq\frac{3}{4}\mathcal{L}_{\mathrm{TE}}^{2}.$
For the proof of Theorem 6.11 see Appendix F.
Boostraping.
A natural question to ask is whether extending the weak-to-strong setup described in Section 2 to use multiple students helps improve over just one student, similar to boosting. We show that allowing multiple students intead of just one would still run into the same limitation of weak-to-strong generalization, under a similar set of conditions on the optimaliy of the teacher and student predictors.
Theorem 6.12 (Limitation of Weak-to-Strong Generlization with a Bounded Student).
If the target $f^{*}$ is normalized, $\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}[(f^{*}({\bm{x}}))^{2}]=1$, then for any teacher $f_{\text{teacher}}$ shrinking-optimal w.r.t. $f^{*}$ and any student $f_{\text{student}}$ has smaller norm than the teacher, i.e., $\|f_{\text{student}}\|_{\mathcal{D}}\leq\|f_{\text{teacher}}\|_{\mathcal{D}}$, we have that
${\mathcal{L}_{\mathrm{ST}}}\geq\left(1-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}\geq\frac{1}{4}\mathcal{L}_{\mathrm{TE}}^{2}$
Corollary 6.13 (Limitation of Weak-to-Strong Generalization with Bootstrapping).
If the target $f^{*}$ is normalized, $\mathbb{E}_{{\bm{x}}\sim\mathcal{D}}[(f^{*}({\bm{x}}))^{2}]=1$, then for any teacher $f_{\text{teacher}}$ shrinking-optimal w.r.t. $f^{*}$ and any chain of student $f^{(i)}_{\text{student}}$, $i=0,2,\ldots,\infty$, where $f^{(i)}_{\text{student}}=f_{\text{teacher}}$ and $f^{(i+1)}_{\text{student}}$ is shrinking-optimal w.r.t. $f^{(i)}_{\text{student}}$, we have that for all $i\geq 1$
$\mathcal{L}_{\mathrm{ST},i}\geq\left(1-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}\geq\frac{1}{4}\mathcal{L}_{\mathrm{TE}}^{2},$
where $\mathcal{L}_{\mathrm{ST},i}$ is the test loss of the $n$-th student. In particular, this include the case where the student $f_{\text{student}}^{(i)}$ is trained by gradient flow w.r.t. $f_{\text{student}}^{(i-1)}$ for arbitrary stopping time $T_{i}$, kernel $\mathcal{K}_{i}$ (potentially induced by different features).
For the proof of Theorem 6.12 and Corollary 6.13 see Appendix F.
7 Conclusions
We prove that weak-to-strong generalization can happen even in a simple model of two-layer neural networks with random features. We show that in this case the PGR (Performance Gap Recovered) can converge to $1$ when teacher loss goes to 00. Additionally, we show an even stronger separation between the teacher and the student, namely that the student error is polynomially smaller than the teacher error. For the ReLU network, we show that the exponent in this polynomial depencence is at least $1.49$ while for the linear networks we show that it is $2$. Further, we explain that early stopping is crucial for weak-to-strong generalization to occur. Most interestingly, we show that there is an inherent limitation to weak-to-strong generalization in random feature models trained with gradient descent or gradient flow and even more generally. Namely, we show that if the target is normalized, the student error can be at most quadratically smaller than the teacher error. Our quadratic lower bound asserts that the student loss cannot be smaller than square of teacher loss, regardless of the choice of student feature, the early stopping time, and even the number of rounds of boostraping.
Acknowledgements
ZL and SA are supported by OpenAI superalignemnt grant. SA is also supported by NSF, DARPA, ONR. MM would like to thank Theodor Misiakiewicz for providing the reference (Defilippis et al., 2024) and for a useful discussion about it.
Appendix
Appendix A Gradient Flow Dynamics
Theorem A.1 (Closed Form Solution to Gradient Flow Dynamics).
Let $\mathcal{K}:\mathcal{X}\times\mathcal{X}\to\mathcal{R}$ be a p.s.d. kernel with eigendecomposition $\mathcal{K}({\bm{x}},{\bm{x}}^{\prime})=\sum_{i=1}^{\infty}\lambda_{i}e_{i}({\bm{x}})e_{i}({\bm{t}})$, where $\langle{e_{i}},{e_{j}}\rangle_{\mathcal{D}}=\delta_{ij}$. For any $f^{*}:\mathcal{X}\to\mathbb{R}$, the equation of gradient flow for a predictor $f_{t}$
[…]
with initial condition given by $f_{0}=0$ has a closed form solution $f_{t}$
$f_{t}=\sum_{k\geq 1}\left(1-e^{-\lambda_{k}t}\right)\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}e_{k}.$
Proof of Theorem A.1.
To solve the gradient flow equation in (16), we project it onto the eigenbasis of the kernel operator $\mathcal{K}$. Let $\beta_{t}^{i}=\langle f_{t},e_{i}\rangle_{\mathcal{D}}$ be the coefficients of $f_{t}$.
Taking the inner product of both sides of (16) with $e_{j}$:
$\frac{d}{dt}\beta_{t}^{j}$ $=-\mathbb{E}_{y\sim\mathcal{D}}\left[\left\langle{\mathcal{K}}(\cdot,y),e_{j}\right\rangle_{\mathcal{D}}(f_{t}(y)-f^{*}(y))\right]$
$=-\lambda_{j}\langle e_{j},f_{t}-f^{*}\rangle_{\mathcal{D}}$
$=-\lambda_{j}(\beta_{t}^{j}-\langle f^{*},e_{j}\rangle_{\mathcal{D}})$
Solving this ODE with initial condition $\beta_{0}^{j}=0$:
$\beta_{t}^{j}=(1-e^{-\lambda_{j}t})\langle f^{*},e_{j}\rangle_{\mathcal{D}}$
Therefore:
$f_{t}=T^{\mathcal{K}}_{t}f^{*}=\sum_{j=1}^{\infty}(1-e^{-\lambda_{j}t})\langle f^{*},e_{j}\rangle_{\mathcal{D}}e_{j}$
Appendix B Different Model Types
We will first introduce a model that is slightly more general than 2.3.
Model B.1 (Diagonal Covariate Features).
Let $\{e_{i}\}_{i=1}^{\infty}$ be a fixed orthonormal basis of $L_{2}^{\mathcal{D}}(\mathcal{X})$. Take the random features to be given by
$g=\sum_{i=1}^{\infty}\langle g,e_{i}\rangle_{\mathcal{D}}e_{i}$
where $\{\langle g,e_{i}\rangle_{\mathcal{D}}\}_{i=1}^{\infty}\sim N(0,\Lambda)$ where $\Lambda=(\lambda_{1},\lambda_{2},\dots)$ in decreasing order.
Our results will hold for B.1 and they are more general than 2.3.
Proposition B.2 (Linear network as Diagonal Coavariate Features).
To get 2.3 from B.1, we take $\Lambda=(\psi_{1},\dots,\psi_{d},0,0,\dots)$, $\langle x\mapsto x^{\top}u,x\mapsto x^{\top}v\rangle_{\mathcal{D}}=\sum_{i=1}^{d}\psi_{i}u_{i}v_{i}$, and $e_{i}(x)=x_{i}/\sqrt{\psi_{i}}$. Thus if we set $g_{i}$ in B.1 as $g(x)={\bm{x}}^{\top}{\bm{u}}_{i}$, where ${\bm{U}}\sim N(0,I_{d})$, then $\{\langle g,e_{i}\rangle_{\mathcal{D}}\}_{i=1}^{\infty}\sim N(0,\Lambda)$.
Proposition B.3 (ReLU Network as Diagonal Covariate Features).
If in B.1 we take $\mathcal{D}=\text{Unif}(\mathbb{S}^{d-1})$, $\{e_{i}\}_{i=1}^{\infty}$ to be the spherical harmonics $\{\{\phi_{i,k}\}_{i=1}^{N_{k}}\}_{k=1}^{\infty}$, and choose $\Lambda=(\sigma_{1},\dots,\sigma_{1},\sigma_{2},\dots,\sigma_{2},\dots,\sigma_{k},\dots,\sigma_{k},\dots)$, where $\sigma_{k}$ repeats $N_{k}$ times, we recover the same feature covariance as in 2.2. Under Gaussian Universality assumptions, for this instance of B.1, the teacher and the student risk behave the same as in 2.2.
Proof of Proposition B.3.
In 2.2, the kernel induce by the random features is given by
${\mathcal{K}}({\bm{x}},{\bm{x}}^{\prime}):=\mathbb{E}_{{\bm{u}}\sim\text{unif}(\mathbb{S}^{d-1})}\left[\sigma(\langle{\bm{u}},{\bm{x}}\rangle)\sigma(\langle{\bm{u}},{\bm{x}}^{\prime}\rangle)\right].$
A random feature in the ReLU case 2.2 is given by
$\sigma(\langle{{\bm{w}}},{{\bm{x}}}\rangle)=\sum_{k\geq 0}N_{k}\sigma_{k}P_{d,k}(\langle{{\bm{w}}},{{\bm{x}}}\rangle)=\sum_{k\geq 0}\sigma_{k}\langle{{\bm{\phi}}_{k}({\bm{w}})},{{\bm{\phi}}_{k}({\bm{x}})}\rangle=\sum_{k=0}^{\infty}\sum_{i=1}^{N_{k}}\sigma_{k}\phi_{i,k}({\bm{w}})\phi_{i,k}({\bm{x}}).$
A random feature in the Diagonal Feature Covariance case B.1 is given by
$g({\bm{x}})=\sum_{k=0}^{\infty}\sum_{i=1}^{N_{k}}g_{i,k}\phi_{i,k}({\bm{x}}).$
$\mathcal{K}({\bm{x}},{\bm{x}}^{\prime})=\mathbb{E}_{{\bm{u}}}[\sigma(\langle{{\bm{u}}},{{\bm{x}}}\rangle)\sigma(\langle{{\bm{u}}},{{\bm{x}}^{\prime}}\rangle)]=\mathbb{E}_{{\bm{u}}}\left[\sum_{k\geq 0}\sigma_{k}\langle{{\bm{\phi}}_{k}({\bm{u}})},{{\bm{\phi}}_{k}({\bm{x}})}\rangle\cdot\sigma_{k}\langle{{\bm{\phi}}_{k}({\bm{u}})},{{\bm{\phi}}_{k}({\bm{x}}^{\prime})}\rangle\right]=\sum_{k\geq 0}\sigma_{k}^{2}\langle{{\bm{\phi}}_{k}({\bm{x}})},{{\bm{\phi}}_{k}({\bm{x}}^{\prime})}\rangle.$
On the other hand, we have that for B.1 the kernel is
$\mathcal{K}({\bm{x}},{\bm{x}}^{\prime})=\mathbb{E}_{g}\left[g({\bm{x}})g({\bm{x}}^{\prime})\right]=\mathbb{E}_{g}\left[\sum_{i,k}\sum_{j,l}g_{i,k}\phi_{i,k}({\bm{x}})g_{j,l}\phi_{j,l}({\bm{x}})\right]=\sum_{k=0}^{\infty}\sigma_{k}^{2}\langle{\bm{\phi}}_{k}({\bm{x}}),{\bm{\phi}}_{k}({\bm{x}}^{\prime})\rangle.$
B.1 Diagonalizability of the Feature Covariance in 2.3
In this section we establish the generality of 2.3. Namely, we show that one can assume that ${\bm{u}}$ is isotropic Gaussuian distribution and ${\bm{x}}$ has diagonal covariance if and only if their covariance matrices are codiagonalizable.
Proposition B.4 (Diagonalizability of the Linear Network Covariance).
A linear network $\sigma(z)=z$ with bottom layer weights ${\bm{u}}$ distributed accoding to a Gaussian distribution $\mathcal{U}=N(0,\Psi_{{\bm{u}}})$ and the inputs ${\bm{x}}$ distributed according to $\mathcal{D}=N(0,\Psi_{{\bm{x}}})$ is equivalent to 2.3 if and only if $\Psi_{{\bm{x}}}$ and $\Psi_{{\bm{u}}}$ are codiagonalizable, with $\Psi={\bm{V}}^{T}{\bm{S}}_{{\bm{u}}}^{-1/2}{\bm{V}}^{T}{\bm{S}}_{{\bm{x}}}{\bm{V}}{\bm{S}}_{{\bm{u}}}^{-1/2}{\bm{V}}$ where ${\bm{V}}$ is the shared basis.
Proof of Proposition B.4.
Note that the distribution in the first case is given by $z\mid{\bm{x}}\sim\langle{\bm{u}},{\bm{x}}\rangle\mid{\bm{x}}\sim N(0,{\bm{x}}\Sigma_{u}{\bm{x}}^{T})$. Note further that for 2.3 with $\Psi$ we have that $\tilde{z}\mid\tilde{\bm{x}}\sim\langle\tilde{\bm{u}},\tilde{\bm{x}}\rangle\mid\tilde{\bm{x}}\sim N(0,\Psi)$. Now if $\tilde{\bm{x}}={\bm{A}}{\bm{x}}$ for some ${\bm{A}}$, then we have that $\Psi=cov(\tilde{\bm{x}})=cov({\bm{A}}{\bm{x}})={\bm{A}}^{T}\Sigma_{x}{\bm{A}}={\bm{A}}^{T}{\bm{V}}_{x}{\bm{S}}_{x}{\bm{V}}_{x}{\bm{A}}$, where ${\bm{V}}_{x}$ is the basis of $\Sigma_{x}$, i.e. $\Sigma_{x}={\bm{V}}_{x}^{T}{\bm{S}}_{x}{\bm{V}}_{x}$ for ${\bm{S}}_{x}$ diagonal. Since $\Psi$ is diagonal. So ${\bm{A}}^{T}{\bm{V}}_{x}{\bm{S}}_{x}{\bm{V}}_{x}{\bm{A}}$ is diagonal if and only if ${\bm{A}}={\bm{D}}{\bm{V}}_{x}$ for ${\bm{V}}_{x}$ a basis of $\Sigma_{x}$ and some diagonal matrix ${\bm{D}}$. So the distributions of $z$ and $\tilde{z}$ are the same if and only if ${\bm{A}}={\bm{D}}{\bm{V}}_{x}$ and ${\bm{x}}\Sigma_{{\bm{u}}}{\bm{x}}^{T}={\bm{x}}{\bm{V}}_{x}{\bm{D}}^{2}{\bm{V}}_{x}{\bm{x}}^{T}$, so $\Sigma_{{\bm{u}}}={\bm{V}}_{x}{\bm{D}}^{2}{\bm{V}}_{x}$. ∎
Appendix C Proof of Weak-to-Strong Multiplicative Error Improvement
Here we present the proof of Theorem 6.3. We will introduce the following notation. Let $M_{\mathrm{TE}}=m$ and let $\{e_{i}\}_{i=1}^{\infty}$ be the eigenbasis of the student’s kernel $\mathcal{K}$. Let the units, i.e. features, of the student be functions $g$ drawn from some distribution $\mathcal{G}$. For example, in 2.2, we gave that $\mathcal{G}$ is given by $g({\bm{x}})=\sigma(\langle{\bm{u}},{\bm{x}}\rangle)$ where ${\bm{u}}\sim\text{unif}(\mathbb{S}^{d-1})$. Let ${\bm{w}}\in\mathbb{R}^{m}$ be the teacher’s weights, and let the teacher be parametrized as $f_{\text{teacher}}=f_{w}=\sum_{i=1}^{m}w_{i}g_{i}$, where $\{g_{i}\}_{i=1}^{m}$ are teacher’s units, i.e. features. We will refer to $\hat{f}_{w}$ the optimally trained teacher. Let the student’s predictor at time $T$ be $f_{T}$. Let $\mathcal{H}$ be the Hilbert space of square-integrable functions with respect to $\mathcal{D}$. Let ${\bm{\Phi}}\in\mathbb{R}^{m\times m}$ be the Gram matrix of random units, i.e., $\Phi_{ij}:=\langle{\sigma_{i}},{\sigma_{j}}\rangle_{\mathcal{D}}=\sum_{k\geq 1}\langle{\sigma_{i}},{e_{k}}\rangle_{\mathcal{D}}\langle{\sigma_{j}},{e_{k}}\rangle_{\mathcal{D}}$.
C.1 Preliminary Lemmas
Lemma C.1.
For all $k,k^{\prime}\geq 1$,
$\mathbb{E}_{g\sim\mathcal{G}}[\langle{g},{e_{k}}\rangle_{\mathcal{D}}\langle{g},{e_{k^{\prime}}}\rangle_{\mathcal{D}}]=\begin{cases}\lambda_{k}&\quad\text{if }k=k^{\prime},\\ 0&\quad\text{otherwise}.\end{cases}$
Lemma C.2.
For $c\in[0,1)$,
$((1-c)(b+\Delta)-b)^{2}\leq\Delta^{2}+\frac{c}{2-c}b^{2}.$
Proof of Lemma C.2.
$((1-c)(b+\Delta)-b)^{2}=((1-c)\Delta-cb)^{2}$ $=\left((1-c)^{2}\cdot\frac{1}{1-c}\Delta+\left(2c-c^{2}\right)\cdot\frac{1}{2c-c^{2}}(-cb)\right)^{2}$
$\leq(1-c)^{2}\cdot\frac{1}{(1-c)^{2}}\Delta^{2}+\left(2c-c^{2}\right)\cdot\frac{b^{2}}{(2-c)^{2}}$
$=\Delta^{2}+\frac{c}{2-c}b^{2},$
where the inequality follows from Jensen’s inequality. ∎
The following is a standard result on the generalized Rayleigh quotient.
Lemma C.3.
If ${\bm{X}},{\bm{Y}}\in\mathbb{R}^{m\times m}$ are symmetric PSD matrices, then
$\sup_{{\bm{u}}\in\mathrm{colspan}({\bm{X}})}\frac{{\bm{u}}^{\top}{\bm{Y}}{\bm{u}}}{{\bm{u}}^{\top}{\bm{X}}{\bm{u}}}=\lambda_{1}\!\left((\sqrt{{\bm{X}}})^{+}{\bm{Y}}(\sqrt{{\bm{X}}})^{+}\right).$
Proof.
Let ${\bm{u}}\in\mathrm{colspan}({\bm{X}})\setminus\{{\bm{0}}\}$. Then ${\bm{u}}=\sqrt{{\bm{X}}}{\bm{z}}$ for some ${\bm{z}}\in\mathbb{R}^{m}\setminus\{{\bm{0}}\}$. Thus,
$\frac{{\bm{u}}^{\top}{\bm{Y}}{\bm{u}}}{{\bm{u}}^{\top}{\bm{X}}{\bm{u}}}=\frac{{\bm{z}}^{\top}(\sqrt{{\bm{X}}})^{+}{\bm{Y}}(\sqrt{{\bm{X}}})^{+}{\bm{z}}}{{\bm{z}}^{\top}{\bm{z}}}\leq\lambda_{1}\!\left((\sqrt{{\bm{X}}})^{+}{\bm{Y}}(\sqrt{{\bm{X}}})^{+}\right).$
The equality is achieved when ${\bm{z}}$ is the top eigenvector of $(\sqrt{{\bm{X}}})^{+}{\bm{Y}}(\sqrt{{\bm{X}}})^{+}$. ∎
C.2 Proof of Theorem 6.3
Lemma C.4.
Under Assumption 6.1, for all $T>0,S\geq K$ and ${\bm{w}}\in\mathbb{R}^{m}$, we have
$\mathcal{L}(f_{T})\leq\mathcal{L}(\hat{f}_{{\bm{w}}})+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\|f^{*}\|_{\mathcal{D}}^{2}-(1-\lambda_{S+1}^{2}T^{2})\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}.$
Proof.
By (13),
$\mathcal{L}(f_{T})$ $=\sum_{k\geq 1}\underbrace{\left(\left(1-e^{-\lambda_{k}T}\right)\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}}_{=:\,\ell_{k}}.$
We provide upper bounds for $\ell_{k}$ in three cases. First, for all $1\leq k\leq K$, by Lemma C.2 we have
$\ell_{k}=\left(\left(1-e^{-\lambda_{k}T}\right)\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}$ $\leq\left(\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}+\frac{e^{-\lambda_{k}T}}{2-e^{-\lambda_{k}T}}\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}^{2}$
$\leq\left(\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}^{2},$
where we used the fact that $e^{-\lambda_{k}T}\leq e^{-\lambda_{K}T}$ for $1\leq k\leq K$.
Then, for $K+1\leq k\leq S$, we have
$\ell_{k}=\left(\left(1-e^{-\lambda_{k}T}\right)\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}=\left(1-e^{-\lambda_{k}T}\right)^{2}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}\leq\left(\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}.$
Finally, for $k\geq S+1$, we have $e^{-\lambda_{k}T}\geq e^{-\lambda_{S+1}T}\geq 1-\lambda_{S+1}T$, and thus $\left(1-e^{-\lambda_{k}T}\right)^{2}\leq\lambda_{S+1}^{2}T^{2}$. This implies the following bound for $\ell_{k}$:
$\ell_{k}=\left(\left(1-e^{-\lambda_{k}T}\right)\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}=(1-e^{-\lambda_{k}T})^{2}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}$ $\leq\lambda_{S+1}^{2}T^{2}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}..$
Putting all these together proves the following:
$\mathcal{L}(f_{T})$ $\leq\sum_{k=1}^{S}\left(\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\sum_{k=1}^{K}\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}^{2}+\lambda_{S+1}^{2}T^{2}\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}$
$=\sum_{k\geq 1}\left(\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}-\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}\right)^{2}+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\sum_{k=1}^{K}\langle{f^{*}},{e_{k}}\rangle_{\mathcal{D}}^{2}-(1-\lambda_{S+1}^{2}T^{2})\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}$
$=\mathcal{L}(\hat{f}_{{\bm{w}}})+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\|f^{*}\|_{\mathcal{D}}^{2}-(1-\lambda_{S+1}^{2}T^{2})\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2},$
which completes the proof. ∎
It remains to bound $\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}$. Let $G:\mathbb{R}^{m}\to{\mathcal{H}}$ be the linear operator that maps ${\bm{u}}\in\mathbb{R}^{m}$ to $\sum_{i=1}^{m}u_{i}g_{i}\in{\mathcal{H}}$. Let $Pf:=\sum_{k=1}^{S}\langle{f},{e_{k}}\rangle e_{k}$ be the projection operator onto the top-$K$ eigenspace of ${\mathcal{K}}$, and $Qf$ be the projection onto the span of $\{g_{1},\dots,g_{m}\}$. Then $G^{*}G={\bm{\Phi}}$, $G^{*}PG={\bm{A}}$ and $Q=G^{*}{\bm{\Phi}}^{+}G$.
In the following, we give a characterization of $\kappa$.
Lemma C.5.
If ${\bm{A}}\neq{\bm{0}}$, then $R:=QPQ\neq 0$ and $\kappa$ equals the smallest non-zero eigenvalue of $R$.
Proof.
If ${\bm{A}}\neq{\bm{0}}$, then there exist $1\leq i<j\leq m$ such that $\langle{Pg_{i}},{Pg_{j}}\rangle\neq 0$. This implies $R\neq 0$ because $\langle{g_{i}},{Rg_{j}}\rangle=\langle{PQg_{i}},{PQg_{j}}\rangle\neq 0$.
Let $\rho_{\min}$ be the smallest non-zero eigenvalue of $R$. To see why $\kappa=\rho_{\min}$, we first express $\rho_{\min}$ as the minimum of a Rayleigh quotient:
$\rho_{\min}=\inf_{\alpha\in\mathrm{range}(R)}\frac{\langle{\alpha},{R\alpha}\rangle}{\langle{\alpha},{\alpha}\rangle}.$ (17)
Then we have
$\mathrm{range}(R)=\mathrm{range}(QPQ)=\{QPG{\bm{v}}\mid{\bm{v}}\in\mathbb{R}^{m}\}$ =$=\{GG^{*}QPG{\bm{v}}\mid{\bm{v}}\in\mathbb{R}^{m}\}$
=$=\{GG^{*}PG{\bm{v}}\mid{\bm{v}}\in\mathbb{R}^{m}\}.$
Since ${\bm{A}}=G^{*}PG$, we further have
$\mathrm{range}(R)=\{G{\bm{A}}{\bm{v}}\mid{\bm{v}}\in\mathbb{R}^{m}\}=\{G{\bm{u}}\mid{\bm{u}}\in\mathrm{colspan}({\bm{A}})\}.$
Plugging this into (17), we have
$\rho_{\min}=\inf_{{\bm{u}}\in\mathrm{colspan}({\bm{A}})}\frac{\langle{G{\bm{u}}},{RG{\bm{u}}}\rangle}{\langle{G{\bm{u}}},{G{\bm{u}}}\rangle}=\inf_{{\bm{u}}\in\mathrm{colspan}({\bm{A}})}\frac{\langle{{\bm{u}}},{G^{*}PG{\bm{u}}}\rangle}{\langle{{\bm{u}}},{G^{*}G{\bm{u}}}\rangle}$ $=\inf_{{\bm{u}}\in\mathrm{colspan}({\bm{A}})}\frac{{\bm{u}}^{\top}{\bm{A}}{\bm{u}}}{{\bm{u}}^{\top}{\bm{\Phi}}{\bm{u}}}$
$=\left(\sup_{{\bm{u}}\in\mathrm{colspan}({\bm{A}})}\frac{{\bm{u}}^{\top}{\bm{\Phi}}{\bm{u}}}{{\bm{u}}^{\top}{\bm{A}}{\bm{u}}}\right)^{-1}.$
Since ${\bm{\Phi}}={\bm{A}}+{\bm{B}}$, we further have
$\rho_{\min}=\left(\sup_{{\bm{u}}\in\mathrm{colspan}({\bm{A}})}\frac{{\bm{u}}^{\top}({\bm{A}}+{\bm{B}}){\bm{u}}}{{\bm{u}}^{\top}{\bm{A}}{\bm{u}}}\right)^{-1}=\left(1+\sup_{{\bm{u}}\in\mathrm{colspan}({\bm{A}})}\frac{{\bm{u}}^{\top}{\bm{B}}{\bm{u}}}{{\bm{u}}^{\top}{\bm{A}}{\bm{u}}}\right)^{-1}.$
By Lemma C.3, the supermum here equals $\lambda_{1}\!\left((\sqrt{{\bm{A}}})^{+}{\bm{B}}(\sqrt{{\bm{A}}})^{+}\right)$. Therefore,
$\rho_{\min}=\frac{1}{1+\lambda_{1}\left((\sqrt{{\bm{A}}})^{+}{\bm{B}}(\sqrt{{\bm{A}}})^{+}\right)}=\kappa,$
which completes the proof. ∎
Now we give a lower bound for $\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}$ in terms of $\mathcal{L}(\hat{f}_{{\bm{w}}})$.
Lemma C.6.
Under Assumptions 6.1 and 6.2, if ${\mathcal{L}}(\hat{f}_{{\bm{w}}})=\min_{{\bm{w}}^{\prime}\in\mathbb{R}^{m}}{\mathcal{L}}(\hat{f}_{{\bm{w}}^{\prime}})$, then
$\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}\geq\kappa\cdot\mathcal{L}(\hat{f}_{{\bm{w}}}),$ (18)
where the equality is attained iff $f^{*}=Ph^{*}$ for some $h^{*}\in{\mathcal{H}}$ that is in the span of the eigenvectors of $R:=QPQ$ corresponding to the eigenvalue $1$ (if exists) and the smallest non-zero eigenvalue, which is $\kappa$.
Proof of Lemma C.6.
By Assumption 6.2, $f^{*}\in\mathrm{span}\{Pg_{1},\dots,Pg_{m}\}=\mathrm{range}(PG)=\mathrm{range}(PQ)$, which implies that there exists $h\in{\mathcal{H}}$ such that $f^{*}=PQh$. As ${\bm{w}}$ minimizes ${\mathcal{L}}(\hat{f}_{{\bm{w}}})$, $\hat{f}_{{\bm{w}}}$ must be the least square solution and $\hat{f}_{{\bm{w}}}=Qf^{*}=QPQh=Rh$. So we have the following identities:
$\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}=\|(I-P)\hat{f}_{{\bm{w}}}\|_{{\mathcal{D}}}^{2}=\|(I-P)QPQh\|_{{\mathcal{D}}}^{2}=\left\langle{h},{QPQ(I-P)QPQh}\right\rangle_{{\mathcal{D}}},$
$\mathcal{L}(\hat{f}_{{\bm{w}}})=\|f^{*}-\hat{f}_{{\bm{w}}}\|_{{\mathcal{D}}}^{2}=\|PQh-QPQh\|^{2}=\left\|(I-Q)PQh\right\|^{2}=\left\langle{h},{QP(I-Q)PQh}\right\rangle_{{\mathcal{D}}}.$
We can express the linear operators $QPQ(I-P)QPQ$ and $QP(I-Q)PQ$ in terms of $R$ as follows.
$QPQ(I-P)QPQ$ $=(QPQ)(QPQ)-(QPQ)(QPQ)(QPQ)=R(R-R^{2}),$
$QP(I-Q)PQ$ $=QPQ-(QPQ)(QPQ)=R-R^{2}.$
Note that $0\preceq R\preceq I$ since $\langle{f},{QPQf}\rangle_{{\mathcal{D}}}=\langle{Qf},{P(Qf)}\rangle_{{\mathcal{D}}}$ and $0\leq\langle{Qf},{P(Qf)}\rangle_{{\mathcal{D}}}\leq\langle{Qf},{Qf}\rangle_{{\mathcal{D}}}=\langle{f},{f}\rangle_{{\mathcal{D}}}$. Let $Rf=\sum_{i=1}^{r}\rho_{i}\langle{v_{i}},{f}\rangle v_{i}$ be an eigendecomposition of $R$, where $01\geq\rho_{1}\geq\rho_{2}\geq\cdots\geq\rho_{r}>$ 0 and $\{v_{1},\dots,v_{r}\}$ are the corresponding eigenvectors.
By Assumption 6.2, ${\bm{A}}\neq{\bm{0}}$. So by Lemma C.5, $R\neq 0$, $r\geq 1$ and $\rho_{r}=\kappa$. Then we have
$\sum_{k\geq S+1}\langle{\hat{f}_{{\bm{w}}}},{e_{k}}\rangle_{\mathcal{D}}^{2}=\left\langle{h},{(R(R-R^{2}))h}\right\rangle_{{\mathcal{D}}}$ $=\sum_{i=1}^{r}\rho_{i}\cdot(\rho_{i}-\rho_{i}^{2})\langle{h},{v_{i}}\rangle_{{\mathcal{D}}}^{2}$
$\geq\sum_{i=1}^{r}\kappa\cdot(\rho_{i}-\rho_{i}^{2})\langle{h},{v_{i}}\rangle_{{\mathcal{D}}}^{2}=\kappa\cdot\left\langle{h},{(R-R^{2})h}\right\rangle_{{\mathcal{D}}}=\kappa\cdot\mathcal{L}(\hat{f}_{{\bm{w}}}),$
which proves (18).
To understand when the equality is attained, we note that the equality holds iff the inequality in the second line is an equality. This means that for all $1\leq i\leq r$, either $\rho_{i}\in\{1,\kappa\}$ or $\langle{h},{v_{i}}\rangle_{{\mathcal{D}}}=0$. Equivalently, $h$ can be expressed as $h=\mu_{1}h_{1}+\mu_{2}h_{2}$ for some $\mu_{1},\mu_{2}\in\mathbb{R}$, where $h_{1}$ is an eigenvector of $R$ corresponding to eigenvalue $1$ and $h_{2}$ is an eigenvector of $R$ corresponding to eigenvalue $\kappa$.
In this case, $f^{*}=PQh=\mu_{1}PQh_{1}+\mu_{2}PQh_{2}$. This can be further simplified to $f^{*}=\mu_{1}Ph_{1}+\mu_{2}Ph_{2}=Ph$ since $h_{1},h_{2}\in\mathrm{range}(R)\subseteq\mathrm{range}(Q)$. ∎
Proof of Theorem 6.3.
Putting Lemma C.4, Lemma C.6 and Lemma C.5 together, we have
$\mathcal{L}(f_{T})$ $\leq\mathcal{L}(\hat{f}_{{\bm{w}}})+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\|f^{*}\|_{\mathcal{D}}^{2}-(1-\lambda_{S+1}^{2}T^{2})\kappa\cdot\mathcal{L}(\hat{f}_{{\bm{w}}})$
$=\left(1-(1-\lambda_{S+1}^{2}T^{2})\kappa\right)\mathcal{L}(\hat{f}_{{\bm{w}}})+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\|f^{*}\|_{\mathcal{D}}^{2},$
which completes the proof. ∎
C.3 General Lower Bound on $\kappa_{S}$
We will bound
$\kappa_{S}:=\frac{1}{1+\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)}.$
in the general case.
Let $F_{A}$ and $F_{B}$ be defined as follows:
$F_{A}$ $:\text{span}(e_{1},\dots,e_{J})\to\mathbb{R}^{m}$
$F_{A}$ $:h\mapsto\begin{pmatrix}\vdots\\ \langle h,g_{j}\rangle_{\mathcal{D}}\\ \vdots\\ \end{pmatrix}$
$F_{B}$ $:\text{span}(e_{J},e_{J+1},\dots)\to\mathbb{R}^{m}$
$F_{B}$ $:h\mapsto\begin{pmatrix}\vdots\\ \langle h,g_{j}\rangle_{\mathcal{D}}\\ \vdots\\ \end{pmatrix}.$
We will denote $F_{A}^{\dagger}$ and $F_{B}^{\dagger}$ their adjoint operators. We can rewrite the largest eigenvalue of $(\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}$ as
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)=\sup_{v\in\text{span}(e_{1},\dots,e_{J})}\frac{v^{T}F_{A}^{\dagger}F_{B}F_{B}^{\dagger}F_{A}v}{v^{T}F_{A}^{\dagger}F_{A}F_{A}^{\dagger}F_{A}v}$
Let $\mathcal{H}$ be a Hilbert space defined as
={linear mappings from $\mathcal{H}=\{\text{linear mappings from }\text{span}(e_{1},\dots,e_{J})\text{ to }\text{span}(e_{J},e_{J+1},\dots)\}$
with the inner product defined as
$\langle f,f^{\prime}\rangle_{\mathcal{H}}=\sum_{j=1}^{J}\langle f(e_{i}),f^{\prime}(e_{i})\rangle_{\mathcal{D}}.$
We will show the following lower bound first.
Proposition C.7.
For $v\in\text{span}(e_{1},\dots,e_{J})$ it holds that
$v^{T}F_{A}^{\dagger}F_{A}F_{A}^{\dagger}F_{A}v\geq\lambda_{J}({\bm{A}})^{2}\|v\|_{2}^{2}.$
Proof.
Let $\sum_{i=1}^{J}u_{i}e_{i}=F_{A}^{\dagger}F_{A}v$. Then we have that
$v^{T}F_{A}^{\dagger}F_{A}F_{A}^{\dagger}F_{A}v\geq\langle\sum_{i=1}^{J}u_{i}e_{i},\sum_{i=1}^{J}u_{i}e_{i}\rangle_{\mathcal{D}}\geq\|u\|_{2}^{2}\geq\lambda_{J}({\bm{A}})^{2}\|v\|_{2}^{2}.$
For the upper bound, we will use the concentration of $\|F_{B}^{\dagger}F_{A}v\|_{\mathcal{D}}^{2}$.
Proposition C.8.
For $v=\sum_{j=1}^{J}v_{i}e_{i}$ it holds that
$\|F_{B}^{\dagger}F_{A}v\|_{\mathcal{D}}^{2}\leq\left(\sum_{j=1}^{J}\|F_{B}^{\dagger}F_{A}e_{i}\|_{\mathcal{D}}^{2}\right)\left(\sum_{j=1}^{J}v_{i}^{2}\right).$
Proof.
This follows immediately from Cauchy-Schwarz inequality. ∎
Proposition C.9.
Let $Y\in\mathcal{H}$ be such that $\langle e_{j},Y(v)\rangle=\langle e_{j},g\rangle_{\mathcal{D}}\langle g,v\rangle_{\mathcal{D}}$
$\|Y\|_{H}^{2}=\sum_{j=1}^{J}\sum_{k=J+1}^{\infty}\langle e_{j},g\rangle_{\mathcal{D}}^{2}\langle g,e_{k}\rangle_{\mathcal{D}}^{2}$
so $\|Y\|_{H}^{2}=\|P_{\leq J}g\|^{2}\|P_{\geq J+1}g\|^{2}$, that is $\|Y\|_{H}=\|P_{\leq J}g\|_{\mathcal{D}}\|P_{\geq J+1}g\|_{\mathcal{D}}$.
Furthermore, let $Y_{i}$ be the r.v. corresponding to the draw of the $i$-th feature $g_{i}$. Then
$F_{B}^{\dagger}F_{A}=\sum_{i=1}^{m}Y_{i}.$
We have the following concentration bound in a Hilbert space.
Theorem C.10 (Pinelis (1994)).
Let $Y_{i}$ be independent $\mathcal{H}$ valued random variable with $\mathbb{E}(Y_{i})=0$, where $\mathcal{H}$ is a seperable Hilbert space. Assume that for every $q\geq 2,q\in\mathbb{N}$ we have that $\mathbb{E}\left(\|Y_{i}\|_{\mathcal{H}}^{q}\right)\leq\frac{1}{2}q!B^{2}L^{q-2}$. Then, for all $n\in\mathbb{N}$ and $\epsilon>0$ we have
$P\left(\|\frac{1}{n}\sum_{i=1}^{n}Y_{i}\|_{\mathcal{H}}>\epsilon\right)\leq 2\exp\left(-\frac{n\epsilon^{2}}{B^{2}+L\epsilon+B\sqrt{B^{2}+2L\epsilon}}\right).$
Putting these together in the bound for $\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)$, we get the following.
Proposition C.11.
The largest eigenvalue of $(\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}$ is bounded by
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq\frac{\sum_{j=1}^{J}\|F_{B}^{+}F_{A}e_{j}\|_{\mathcal{D}}^{2}}{\lambda_{J}({\bm{A}})^{2}}$
Proof of Proposition C.11.
This follows directly from Proposition C.8 and Proposition C.7. ∎
We can now use this to bound the norm of $F_{B}^{\dagger}F_{A}$.
Proposition C.12 (Hilbert Space Concentration).
With probability at least $1-2\delta$
$\|F_{B}^{\dagger}F_{A}\|_{\mathcal{H}}^{2}\leq mB\left(\log\frac{1}{\delta}\right)^{2}$
where $B$ is as defined in Theorem C.10.
Proof of Proposition C.12.
In Theorem C.10, taking $m^{2}{\epsilon}^{2}=mB^{2}\log^{2}\frac{1}{\delta}$, we get the desired result. ∎
We can lower bound the eigenvalues of ${\bm{A}}$ using matrix concentration results.
Theorem C.13.
Let ${\bm{a}}=\begin{pmatrix}\vdots\\ \langle g,e_{k}\rangle\\ \vdots\end{pmatrix}_{k=1}^{J}$ as $g\sim\mathcal{G}$ be a row of ${\bm{A}}$. Depending on whether the rows of ${\bm{A}}$ are subgaussian or bounded a.s. we have the following two bounds:
  1. 1.
    Under B.1, i.e. with $D_{g}=N(0,\Lambda)$ where $\Lambda=\text{diag}(\lambda_{1},\lambda_{2},\dots)$, then with probability $1-2\exp(-ct_{A}^{2})$ it holds that
    $\lambda_{J}({\bm{A}})\geq\lambda_{J}\left(\sqrt{m}-C\sqrt{J}-t_{A}\right)^{2},$
    if $\sqrt{m}\geq C\sqrt{J}+t_{A}$, where $C=\Theta(L_{g}^{2})$ and $c=\Theta(1/L_{g}^{4})$ where $L_{g}$ is the subgaussian norm of the vector ${\bm{a}}$. In this case $L_{g}\leq C_{g}\lambda_{1}$, where $C_{g}$ is an absolute constant.
  2. 2.
    Under 2.2, i.e. if $g=\sigma(\langle w,x\rangle)$ with $w\sim\text{unif}(S^{d-1})$ and $N_{0}+N_{1}+\dots+N_{s-1}+1\leq J\leq N_{0}+N_{1}+\dots+N_{s}$ and $s$ is $1$ or even, we have that with probability $1-2J\exp(-ct_{A}^{2})$
    $\lambda_{J}({\bm{A}})\geq\sigma_{s}^{2}\left(\sqrt{m}-t_{A}\sqrt{J}\right)^{2},$
    if $\sqrt{m}\geq t_{A}\sqrt{J}$, where $c$ is an absolute constant.
Under 6.4, with probability $1-2J\exp(-ct_{A}^{2})$ we have that
$\lambda_{J}({\bm{A}})\geq\sigma_{s}^{2}\left(\sqrt{m}-t_{A}\sqrt{J}\right)^{2},$
This computation allows us to lower bound $\kappa_{S}$ in the cases of 2.2 in Appendix E and 2.3 in B.1.
Proof of Theorem C.13.
We will use Theorems $5.39$ and $5.41$ in (Vershynin, 2011). These theorems say that if the rows of a matrix are either bounded a.s. or sub-gaussian, then the singular values of this matrix are upper and lower bounded. Note first that in both cases, ${\bm{A}}$ is a PSD matrix, so it’s sufficient to consider the singular values of ${\bm{A}}$. In both cases, note that in order to apply the theorems from (Vershynin, 2011) to a matrix ${\bm{X}}$, we need to have $\Sigma({\bm{X}})={\bm{I}}$, i.e. the rows of ${\bm{X}}$ have to be isotropic random variables. Note that in both cases
$\Sigma({\bm{A}})_{ij}=\mathbb{E}_{\mathcal{G}}\left(\langle g,e_{i}\rangle\langle g,e_{j}\rangle\right).$
If we let $\Sigma=\Sigma({\bm{A}})$, we can apply these theorems to $\Sigma^{-\frac{1}{2}}{\bm{G}}$, where ${\bm{G}}$ is such that ${\bm{A}}={\bm{G}}^{T}{\bm{G}}$. In that case, the theorems say that with corresponding probabilities that
$s_{J}(\Sigma^{-\frac{1}{2}}{\bm{G}})$ $\geq(\sqrt{m}-c\sqrt{J}-t_{A})^{2}$
$s_{J}(\Sigma^{-\frac{1}{2}}{\bm{G}})$ $\geq(\sqrt{m}-t_{A}\sqrt{J})^{2},$
respectively, where $s_{J}$ is the $J$-th singular value of a matrix. Note that we have that
$s_{J}({\bm{G}})$ $\geq s_{\min}(\Sigma^{\frac{1}{2}})s_{J}(\Sigma^{-\frac{1}{2}}{\bm{G}}).$
Therefore, we get that with corresponding probabilities
$\lambda_{J}({\bm{A}})$ $\geq\lambda_{J}(\Sigma)\left(s_{J}(\Sigma^{-\frac{1}{2}}{\bm{G}})\right)^{2}.$
Here, we need lower bounds on $m$ so that we can square the inequality and go from singular values to eigenvalues.
In the case of 2.3, we have that
$\Sigma({\bm{A}})_{ij}=\Lambda.$
Therefore, we have that $\lambda_{J}(\Sigma)=\lambda_{J}$. Note that we are only left with verifying the conditions of the theorem. First is that the rows of $\Sigma^{-\frac{1}{2}}{\bm{A}}$ are isotropic which is true by design. Second, we need that the rows of $\Sigma^{-\frac{1}{2}}{\bm{A}}$ are subgaussian with a subgaussian constant $L_{g}$. Note note that if $X\sim N(0,\sigma)$ then $\|X\|_{\psi_{2}}\leq C\sigma$ for some absolute constant $C$. Note also that $\|{\bm{a}}\|_{\psi_{2}}=\sup_{x\in S^{J-1}}\|x^{T}{\bm{a}}\|_{\psi_{2}}\leq\sqrt{C_{1}\sum_{i=1}^{J}x_{i}^{2}\|{\bm{a}}_{i}^{2}\|_{\psi_{2}}^{2}}\leq\sqrt{C_{1}C\sum_{i=1}^{J}x_{i}^{2}\lambda_{i}^{2}}\leq\sqrt{C_{1}C\lambda_{1}^{2}\sum_{i=1}^{J}x_{i}^{2}}=\sqrt{C_{1}C\lambda_{1}^{2}}=C_{g}\lambda_{1}$. Note that therefore under 2.3, the conditions of the Theorem $5.39$ in Vershynin (2011) are satisfied, so we get that with probability $1-2\exp(-ct_{A}^{2})$
$\lambda_{J}({\bm{A}})\geq\lambda_{J}\left(\sqrt{m}-C\sqrt{J}-t_{A}\right)^{2},$
as long as $\sqrt{m}\geq c\sqrt{J}+t_{A}$. Here $C=\Theta(L_{g}^{2})$ and $c=\Theta(\frac{1}{L_{g}^{4}})$ in the sense that both $C$ and $c$ are upper and lower bounded by an absolute constant times the appropriate expression of $L_{g}$.
In the case of 2.2, note that since the target Assumption 6.6 will not have anything in the directions with zero $\sigma_{k}$, we can just ignore and reindex the basis functions. So, for now we will just write as if they are nonzero, but we will keep in mind that we skip the zero $\sigma_{k}$. Therefore, we have that
$\Sigma({\bm{A}})_{i,j}=\text{diag}(\sigma_{1}^{2},\sigma_{1}^{2},\dots,\sigma_{2}^{2},\dots,\sigma_{4}^{2},\dots),$
where each $\sigma_{i}$ repeats with the multiplicity $N_{i}$ for $i=1$ or $i$ even (otherwise it’s 00). Furthermore, note that $\|{\bm{a}}\|_{2}^{2}=J$, so we can indeed apply Theorem $5.41$ from (Vershynin, 2011). Therefore, we have that with probability at least $1-2J\exp(-ct_{A}^{2})$
$\lambda_{J}({\bm{A}})\geq\sigma_{s}^{2}(\sqrt{m}-t_{A}\sqrt{J})^{2}$
where $c$ is an absolute constant. ∎
Appendix D Weak-to-Strong Bound Applied to the Case of Linear Networks
Proof of Theorem 3.2.
Follows from Theorem D.5. ∎
Proof of Theorem 6.8.
This follows directly from Proposition B.2 and the general bound Theorem D.1. ∎
Theorem D.1 (Weak-to-Strong Generalization with Gaussian Feature Distribution).
Under Assumption 6.1, if B.1 holds for some $K$, and if the teacher attains minimum loss $\mathcal{L}_{\mathrm{TE}}=\hat{\mathcal{L}}_{\min}$, then for the student trained until time $T$ we have with probability at least $1-\frac{4}{m}$ that
${\mathcal{L}_{\mathrm{ST}}}$ $\leq\inf_{S\geq K}\Biggl{\{}\frac{\bigl{(}\sum_{i=1}^{S}\lambda_{i}\bigr{)}\bigl{(}\sum_{i=S+1}^{\infty}\lambda_{i}\bigr{)}}{\lambda_{S}^{2}}\,\frac{(\log m)^{2}}{m}\,\mathcal{L}_{\mathrm{TE}}+\lambda_{S+1}\,T^{2}\,\mathcal{L}_{\mathrm{TE}}\Biggr{\}}+\frac{e^{-\lambda_{K}T}}{2-e^{-\lambda_{K}T}}\,\bigl{\|}f^{*}\bigr{\|}_{\mathcal{D}}^{2}.$
Proof of Theorem D.1.
This follows from Theorem 6.3 using Theorem D.4 and noting that we can replace $\lambda_{S+1}T^{2}(1-\kappa_{S})\leq\lambda_{S+1}T^{2}$. ∎
D.1 Proofs Related to The Case of Linear Networks
We first aim to show that the $k$-th moments of $\|Y\|_{\mathcal{H}}=\|P_{\leq J}g\|_{\mathcal{D}}\|P_{\geq J+1}g\|_{\mathcal{D}}$ are bounded.
Proposition D.2.
Let $X=\sum_{i=1}^{n}a_{i}^{2}X_{i}^{2}$ where $X_{i}$ are i.i.d. standard Gaussian. Then for $k\geq 2$
$\mathbb{E}\left(X^{k}\right)\leq k!\left(\sum_{i=1}^{m}a_{i}^{2}\right)^{k}.$
Proof.
Expanding the expression for $X^{k}=\left(\sum_{i=1}^{n}a_{i}^{2}X_{i}^{2}\right)^{k}$ and using linearity of expectation, we have that
$\mathbb{E}\left(X^{k}\right)$ $=\sum_{|\alpha|=k}\frac{k!}{\alpha!}\prod_{i=1}^{m}\left((a_{i})^{2}\right)^{\alpha_{i}}\mathbb{E}\left(\left(X_{i}^{2}\right)^{\alpha_{i}}\right).$
Note that $\mathbb{E}(X_{i}^{2\alpha_{i}})=(2\alpha_{i}-1)!!\leq 2^{\alpha_{i}}(\alpha_{i})!$, so we have that
$\mathbb{E}\left(X^{k}\right)$ $=\sum_{|\alpha|=k}\frac{k!}{\alpha!}\prod_{i=1}^{m}\left((a_{i})^{2}\right)^{\alpha_{i}}\mathbb{E}\left(\left(X_{i}^{2}\right)^{\alpha_{i}}\right)\leq\sum_{|\alpha|=k}\frac{k!}{\alpha!}\prod_{i=1}^{m}\left((a_{i})^{2}\right)^{\alpha_{i}}2^{\alpha_{i}}(\alpha_{i})!$
$=\sum_{|\alpha|=k}\frac{k!}{\alpha!}\alpha!\prod_{i=1}^{m}\left(2(a_{i})^{2}\right)^{\alpha_{i}}=k!\sum_{|\alpha|=k}\frac{\alpha!}{\alpha!}\prod_{i=1}^{m}\left(2(a_{i})^{2}\right)^{\alpha_{i}}$
$\leq k!\sum_{|\alpha|=k}\frac{k!}{\alpha!}\prod_{i=1}^{m}\left(2(a_{i})^{2}\right)^{\alpha_{i}}$
where the last inequality is true because $\alpha!\leq k!$. Note that $\sum_{|\alpha|=k}\frac{k!}{\alpha!}\prod_{i=1}^{m}\left(2(a_{i})^{2}\right)^{\alpha_{i}}=\left(\sum_{i=1}^{m}2a_{i}^{2}\right)^{k}$, so the final bound is
$\mathbb{E}\left(X^{k}\right)$ $\leq k!\sum_{|\alpha|=k}\frac{k!}{\alpha!}\prod_{i=1}^{m}\left(2(a_{i})^{2}\right)^{\alpha_{i}}=k!\left(\sum_{i=1}^{m}2a_{i}^{2}\right)^{k}.$
This implies that $\|Y\|_{\mathcal{H}}^{k}$ is bounded for all $k\geq 2$ with an expression of the required form.
Proposition D.3.
Let $g\sim\mathcal{G}$ with $\langle g,e_{i}\rangle\sim N(0,\Lambda)$ with $\Lambda=\text{diag}(\lambda_{1},\lambda_{2},\dots)$. Then for $Y$ it holds that for all $k\geq 2$
$\mathbb{E}\left(\|Y\|_{\mathcal{H}}^{k}\right)\leq k!\left(\sqrt{\left(2\sum_{i=1}^{J}\lambda_{i}\right)\left(2\sum_{i=J+1}^{\infty}\lambda_{i}\right)}\right)^{k}.$
Therefore, in this case we can take $B=\sqrt{2\left(2\sum_{i=1}^{J}\lambda_{i}\right)\left(2\sum_{i=J+1}^{\infty}\lambda_{i}\right)}$ and $L=\sqrt{\left(2\sum_{i=1}^{J}\lambda_{i}\right)\left(2\sum_{i=J+1}^{\infty}\lambda_{i}\right)}$.
Note that Proposition C.12 applies in this case as well. Additionally, Theorem C.13 applies in this case.
Theorem D.4.
Let $g\sim\mathcal{G}$ with $\langle g,e_{i}\rangle\sim N(0,\Lambda)$ with $\Lambda=\text{diag}(\lambda_{1},\lambda_{2},\dots)$. Then with probability at least $1-2\delta-2\exp(-c_{1}\frac{1}{\lambda_{1}^{4}}t_{A}^{2})$ we have that
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq\frac{\|F_{B}^{\dagger}F_{A}\|_{\mathcal{H}}^{2}}{\lambda_{S}({\bm{A}})^{2}}\leq\frac{8m\left(\sum_{i=1}^{S}\lambda_{i}\right)\left(\sum_{i=S+1}^{\infty}\lambda_{i}\right)\left(\log\frac{1}{\delta}\right)^{2}}{\lambda_{S}^{2}\left(\sqrt{m}-C\sqrt{S}-t_{A}\right)^{4}}$
where $c_{1}$ and $C$ are absolute constants. Setting $t_{A}=\lambda_{1}^{2}\log\frac{1}{\delta_{1}}$, $\delta=\frac{1}{m},\delta_{1}=\frac{1}{m}$, we get that with probability at least $1-\frac{4}{m}$ we have
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq\frac{8m\left(\sum_{i=1}^{S}\lambda_{i}\right)\left(\sum_{i=S+1}^{\infty}\lambda_{i}\right)\left(\log m\right)^{2}}{\lambda_{S}^{2}\left(\sqrt{m}-C\sqrt{S}-\lambda_{1}^{2}\log m\right)^{4}}.$
If additionally $m=\omega(S)$ and $m=\omega((\lambda_{1}\log m)^{2})$ then with probaility $1-\frac{4}{m}$
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq 8\frac{\left(\sum_{i=1}^{S}\lambda_{i}\right)\left(\sum_{i=S+1}^{\infty}\lambda_{i}\right)}{\lambda_{S}^{2}}\frac{\left(\log m\right)^{2}}{m}.$
Proof of Theorem D.4.
This theorems follows directly by taking the approproate lower bound from Theorem C.13 and combining that with the Hilbert concentration bound Proposition C.12. ∎
D.2 Additional Proofs for Linear Networks
Proof of Theorem 3.3.
This follows from Theorem D.6 by taking $\epsilon_{d}=0.01$, $\epsilon_{m}=0.01$, $m\geq\Theta(\alpha^{\frac{2}{1-\epsilon_{d}}})$ and $\delta=0.01$. Note that since $m=\sqrt{\frac{d-1}{\alpha}}$ we can turn this inequality into an inequality with $\alpha$ and $d$, i.e $\alpha\geq\Theta(\frac{1}{d^{0.49}})$. ∎
Theorem D.5 (Weak-to-Strong Generalization with Linear Network).
Consider 2.3 when $f^{*}$ is supported by the first $k$ coordinates, i.e. $f^{*}=\beta{\bm{x}}$, where $\beta_{i}=0$ for $i>k$ and take $f^{*}$ to be norm $1$. Take $\Psi_{m}=(1,\dots,1,\frac{1}{m},\dots,\frac{1}{m})$, where $1$ is repeated $k$ times and $\frac{1}{m}$ is repeated $N_{1}(m)=\alpha m^{2}$ times, with $\alpha=m^{-\eta}$ for any $\eta\in[0,\frac{1}{2}]$ (so $d=k+N_{1}(m)$). Then, for all $m>C\sqrt{k}$ 666here $C$ is an absolute constant given by Theorem D.4, if the student is trained until time $T=\log m+\log\frac{1}{\mathcal{L}_{\mathrm{TE}}}$ with probability at least $0.99$,
$\mathcal{L}_{\mathrm{TE}}$ $\geq\frac{\frac{1}{k}m^{-\eta}}{1+\frac{1}{k}m^{-\eta}}$
${\mathcal{L}_{\mathrm{ST}}}$ $\leq k\frac{\log^{2}m}{m^{\eta}}\mathcal{L}_{\mathrm{TE}}+\frac{1}{m}(\log m+\log\frac{1}{\mathcal{L}_{\mathrm{TE}}})^{2}\mathcal{L}_{\mathrm{TE}}+\frac{1}{m}\mathcal{L}_{\mathrm{TE}}$
So, for all $m$ we have with probability at least $0.99$,
${\mathcal{L}_{\mathrm{ST}}}\leq\tilde{O}(k\mathcal{L}_{\mathrm{TE}}^{2}).$
In particular $\mathrm{PGR}\to 1$ as $M_{\mathrm{TE}}\to\infty$.
Proof of Theorem D.5.
For $k=1$ this follows from Theorem D.6. Consider $k$ copies of Theorem D.6 with $N(m)=\frac{\alpha}{k}m^{2}$ each, $\Psi_{m,1},\dots,\Psi_{m,k}$. For example, $\Psi_{m,1}=(1,0,\dots,0,\frac{1}{m},\dots,\frac{1}{m},0,\dots,0)$, were we have $N(m)=\frac{N_{1}(m)}{k}$ repeating $\frac{1}{m}$. Note that a set of weights ${\bm{w}}\in\mathbb{R}^{m}$ that the teacher chooses to optimize over all $k$ problems is worse than choosing a separate set of weights for each of the problems, so the loss will be greater than the sum of the losses in each of the subproblems. So if $\mathcal{L}_{\mathrm{TE}}^{i}$ is the loss in the $i$-th instance, we have that $\mathcal{L}_{\mathrm{TE}}\geq\sum_{i}\mathcal{L}_{\mathrm{TE}}^{i}$. Let $B_{j}^{2}$ be the norm of $f^{*}$ in the directions of the basis functions corrresponding to the $j$-th subset of $\Psi_{m}$, i.e. $B_{j}^{2}=\sum_{i\in\Psi_{m,j}}\beta_{i}^{2}$. Note that by the normalization of $f^{*}$, $\sum_{j=1}^{k}B_{j}^{2}=1$. Furthermore, note that our setup is scale homogeneous, i.e. if we scale $\|f^{*}\|_{\mathcal{D}}^{2}$ by $c^{2}$ then the loss will scale by $c^{2}$ as well. Therefore, we can apply Theorem D.6 to each of the $k$ subprolems to get with probability at least $1-2\exp(-\frac{1}{4}(\alpha m^{2})^{\epsilon_{d}})-2\exp(-\frac{1}{4}m^{\epsilon_{m}})$ that
$\mathcal{L}_{\mathrm{TE}}^{j}\geq B_{j}^{2}\left(0.99\frac{\alpha}{k}\frac{1}{\frac{\alpha}{k}+1}-2m^{-1/2+\epsilon_{m}}-(\alpha m^{2})^{-1/2+\epsilon_{d}}\alpha\right)\geq 0.98B_{j}^{2}\frac{\frac{m^{-\eta}}{k}}{\frac{m^{-\eta}}{k}+1}.$
Therefore, we have that $\mathcal{L}_{\mathrm{TE}}\geq\sum_{j=1}^{k}0.98B_{j}^{2}\frac{\frac{m^{-\eta}}{k}}{\frac{m^{-\eta}}{k}+1}=0.98\frac{\frac{m^{-\eta}}{k}}{\frac{m^{-\eta}}{k}+1}$.
Now consider Theorem D.1 in this case. Set $S=k$. Then we get for $\delta_{T}=\frac{1}{m}\mathcal{L}_{\mathrm{TE}}$ the upper bound since the sums of $\psi_{i}$ are $\sum_{i=1}^{k}\psi_{i}=k$ and $\sum_{i=k+1}^{d}\psi_{i}=\alpha m$ respectively with probability at least $1-\frac{1}{100}\frac{1}{m^{4}}-2\exp(-c_{1}m^{1/4})$
${\mathcal{L}_{\mathrm{ST}}}$ $\leq 40k\alpha\log^{2}m\mathcal{L}_{\mathrm{TE}}+\frac{1}{m}(\log m+\log\frac{1}{\mathcal{L}_{\mathrm{TE}}})^{2}\mathcal{L}_{\mathrm{TE}}+\frac{1}{m}\mathcal{L}_{\mathrm{TE}}$
where the constant $40$ comes from the constant $8$ from the upper bound and the constant from $\log\frac{1}{\delta}$ by setting $\delta=100m^{4}$. Note now that since $\mathcal{L}_{\mathrm{TE}}$ is lower bounded by $\frac{m^{-\eta}}{2k}$, we can combine $\log\frac{1}{\mathcal{L}_{\mathrm{TE}}}$ and $\log m$ terms into
${\mathcal{L}_{\mathrm{ST}}}\leq 50k\frac{\log^{2}m}{m^{\eta}}\mathcal{L}_{\mathrm{TE}}.$
Here we bounded $(\log m+\log 2k)^{2}\leq 2\log m$. This exactly gives with probability at least $1-\frac{1}{10}\frac{1}{m^{4}}-2\exp(-c_{1}m^{1/4})-2\exp(-\frac{1}{4}(\alpha m^{2})^{\epsilon_{d}})-2\exp(-\frac{1}{4}m^{\epsilon_{m}})$
${\mathcal{L}_{\mathrm{ST}}}\leq\tilde{O}(k\mathcal{L}_{\mathrm{TE}}^{2}).$
By setting $\epsilon_{d}<\frac{1}{8}$ and $\epsilon_{m}<\frac{1}{8}$, we get that by union bound over all $m$, with probability at least $0.99$ for all $m$
${\mathcal{L}_{\mathrm{ST}}}\leq\tilde{O}(k\mathcal{L}_{\mathrm{TE}}^{2}).$
In Theorem D.6 we will establish a case where we can provably lower bound the teacher loss.
Theorem D.6 (Lower bound for Teacher Loss).
In Assumption 6.1 take $f^{*}=e_{1}$ and in 2.3 take $\langle g,e_{i}\rangle\sim N(0,\Lambda_{m})$, where $\Lambda_{m}=\text{diag}(1,\frac{1}{m},\frac{1}{m},\dots,\frac{1}{m})$, where $\frac{1}{m}$ is repeated $N(m)$ times. If $N(m)=\alpha m^{2}$ then for all $\epsilon_{d},\epsilon_{m}\in[0,\frac{1}{2}]$ with probability $1-2\exp(-\frac{1}{4}(\alpha m)^{\epsilon_{d}})-2\exp(-\frac{1}{4}m^{\epsilon_{m}})$ we have $\mathcal{L}_{\mathrm{TE}}\geq 0.99\frac{\alpha}{1+\alpha}-2m^{-\frac{1}{2}+\epsilon_{m}}-(\alpha m)^{-\frac{1}{2}+\epsilon_{d}}\alpha$. Furthermore, as long as $\alpha m\left(\log\frac{1}{\delta}\right)^{2}=\Omega(m^{-2+\epsilon_{s}})$ for some $\epsilon_{s}>0$, $1-\delta-2\exp(-c_{1}m^{\frac{1}{4}})$, we have that ${\mathcal{L}_{\mathrm{ST}}}\leq\alpha m(\log(\frac{1}{\delta}))^{2}\mathcal{L}_{\mathrm{TE}}\left(1+O(m^{-\epsilon_{s}})\right)$.
Lemma D.7.
Let ${\bm{G}}_{ij}=\sum_{i=2}^{\infty}\langle g_{j},e_{i}\rangle\langle g_{l},e_{i}\rangle$, let ${\bm{H}}_{i,k}=\langle g_{i},e_{k}\rangle\in\mathbb{R}^{m\times d}$, and let ${\bm{g}}=\begin{pmatrix}\vdots\\ \langle g_{j},e_{1}\rangle\\ \vdots\end{pmatrix}_{j=1}^{m}\sim N(0,\psi{\bm{I}}_{m})$ be the rows of ${\bm{H}}$, i.e. $\Lambda=\psi{\bm{I}}_{m}$. Then with probability at least $1-2\exp(-\frac{d\delta^{2}}{4})$ for all ${\bm{v}}\in\mathbb{R}^{m}$
${\bm{v}}^{T}{\bm{G}}{\bm{v}}\geq(1-\delta)tr(\Lambda){\bm{v}}^{T}{\bm{v}}.$
Proof of Lemma D.7.
Note that ${\bm{G}}={\bm{H}}{\bm{H}}^{T}$ so we have that ${\bm{v}}^{T}{\bm{G}}{\bm{v}}=\|{\bm{H}}^{T}{\bm{v}}||_{2}^{2}$. Not further that
${\bm{H}}^{T}{\bm{v}}=\begin{pmatrix}\vdots\\ \sum_{i=1}^{m}v_{i}\langle g_{i},e_{j}\rangle\\ \vdots\end{pmatrix}_{j=1}^{d}\sim N(0,\psi{\bm{I}}_{d}\|v\|_{2}^{2}),$
since the entries $\sum_{i=1}^{m}v_{i}\langle g_{i},e_{j}\rangle$ are independent and each has variance $\psi\|v\|_{2}^{2}$.
Note now that
$\|{\bm{H}}^{T}{\bm{v}}\|_{2}^{2}=\sum_{j=1}^{d}\left(\sum_{i=1}^{m}v_{i}\langle g_{i},e_{j}\rangle\right)^{2}\sim\psi\|{\bm{v}}\|_{2}^{2}\left(Z_{1}^{2}+\dots+Z_{d}^{2}\right),$
where $Z_{i}\sim N(0,1)$ are iid, i.e. $\|{\bm{H}}^{T}{\bm{v}}\|_{2}^{2}\sim\psi\|{\bm{v}}\|_{2}^{2}X$ where $X\sim\chi^{2}_{d}$. By concentation of $\chi_{d}^{2}$ distribution, we have that with probability at least $1-2\exp(-\frac{d\delta^{2}}{4})$ that
${\bm{v}}^{T}{\bm{G}}{\bm{v}}=\|{\bm{H}}^{T}{\bm{v}}\|_{2}^{2}\geq(1-\delta)\psi d{\bm{v}}^{T}{\bm{v}}=(1-\delta)tr(\Lambda){\bm{v}}^{T}{\bm{v}}.$
Proof of Theorem D.6.
Let $\hat{f}_{\hat{\bm{w}}}$ be the teacher predictor. Note that we can rewrite this as
$\mathbb{E}_{\mathcal{D}}\left(f^{*}-\hat{f}_{{\bm{w}}}\right)^{2}$ $=\mathbb{E}_{\mathcal{D}}\left(e_{1}(x)^{2}\left(1-\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle\right)^{2}\right)+\mathbb{E}_{\mathcal{D}}\left(\left(\sum_{j=1}^{m}w_{j}\sum_{i=2}^{\infty}\langle g_{j},e_{i}\rangle e_{i}(x)\right)^{2}\right)$
$=1-2\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle+\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle\right)^{2}+\mathbb{E}_{\mathcal{D}}\left(\left(\sum_{i=1}^{\infty}\sum_{j=1}^{m}w_{j}\langle g_{j},e_{i}\rangle e_{i}(x)\right)^{2}\right)$
$=1-2\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle+\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle\right)^{2}+\sum_{i=2}^{\infty}\mathbb{E}_{\mathcal{D}}\left(\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{i}\rangle\right)^{2}\right)$
$=1-2\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle+\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle\right)^{2}+\sum_{i=2}^{\infty}\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{i}\rangle\right)^{2}$
Let $[{\bm{G}}]_{jl}=\sum_{i=2}^{\infty}\langle g_{j},e_{i}\rangle\langle g_{l},e_{i}\rangle$ and ${\bm{g}}=\begin{pmatrix}\vdots\\ \langle g_{j},e_{1}\rangle\\ \vdots\end{pmatrix}_{j=1}^{m}$.
Now, we can rewrite the above equation as follows.
$\mathcal{L}(\hat{f}_{{\bm{w}}})$ $=1-2\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle+\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{1}\rangle\right)^{2}+\sum_{i=2}^{\infty}\left(\sum_{j=1}^{m}w_{j}\langle g_{j},e_{i}\rangle\right)^{2}$
$=1-2{\bm{w}}^{T}{\bm{g}}+({\bm{w}}^{T}{\bm{g}})^{2}+{\bm{w}}^{T}{\bm{G}}{\bm{w}}.$
By Lemma D.7, we have that with probability at least $1-2\exp(-\frac{d\delta^{2}}{4})$ that
${\bm{w}}^{T}{\bm{G}}{\bm{w}}\geq(1-\delta)tr(\Lambda){\bm{w}}^{T}{\bm{w}},$
where $d=N(m)$. Therefore, we have that
$\mathcal{L}(\hat{f}_{{\bm{w}}})$ $=1-2{\bm{w}}^{T}{\bm{g}}+({\bm{w}}^{T}{\bm{g}})^{2}+{\bm{w}}^{T}{\bm{G}}{\bm{w}}\geq 1-2{\bm{w}}^{T}{\bm{g}}+({\bm{w}}^{T}{\bm{g}})^{2}+(1-\delta)tr(\Lambda){\bm{w}}^{T}{\bm{w}}.$
The last expression is minimized for ${\bm{w}}=\frac{1}{\|{\bm{g}}\|_{2}^{2}+(1-\delta)tr(\Lambda)}{\bm{g}}$. Note that we have $tr(\Lambda)=\alpha m$. Note further that $\|{\bm{g}}\|_{2}^{2}\sim\chi_{m}^{2}$, so it concetrates around $m$, i.e. with probability at least $1-2\exp(-\frac{\delta_{m}^{2}m}{4})$ we have
$(1+\delta_{m})m\geq\|{\bm{g}}\|_{2}^{2}\geq(1-\delta_{m})m.$
So overall, we have that with probability at least $1-2\exp(-\frac{d\delta^{2}}{4})-2\exp(-\frac{m\delta_{m}^{2}}{4})$
$\mathcal{L}(\hat{f}_{{\bm{w}}})$ $\geq\left(1-\frac{\|{\bm{g}}\|_{2}^{2}}{\|{\bm{g}}\|_{2}^{2}+(1-\delta)tr(\Lambda)}\right)^{2}+(1-\delta)tr(\Lambda)\left(\frac{1}{\|{\bm{g}}\|_{2}^{2}+(1-\delta)tr(\Lambda)}\right)^{2}\|{\bm{g}}\|_{2}^{2}$
$\geq\left(1-\frac{m(1+\delta_{m})}{m(1+\delta_{m})+(1-\delta)\alpha m}\right)^{2}+(1-\delta)\alpha\left(\frac{1}{(1+\delta_{m})+(1-\delta)\alpha}\right)^{2}(1-\delta_{m})$
$=\frac{(1-\delta)\alpha}{1+\delta_{m}+(1-\delta)\alpha}-\frac{2\delta_{m}}{1+\delta_{m}+(1-\delta)\alpha}.$
Theorem D.4 in this case says that for $J=1$ we have with probability $1-\delta-2\exp(-c_{1}m^{\frac{1}{4}})$
$1-\kappa\leq 8\frac{1\cdot\frac{1}{m}N(m)}{m}(\log(\frac{1}{\delta}))^{2}=8\alpha m(\log(\frac{1}{\delta}))^{2}.$
Therefore Theorem 6.3 implies that with probability at least $1-\delta-2\exp(-c_{1}m^{\frac{1}{4}})$ with $T=\log(\frac{1}{\delta_{T}})$
${\mathcal{L}_{\mathrm{ST}}}\leq 8\alpha m(\log(\frac{1}{\delta}))^{2}\mathcal{L}_{\mathrm{TE}}+\frac{1}{m^{2}}(\log(\frac{1}{\delta_{T}}))^{2}\mathcal{L}_{\mathrm{TE}}+\frac{\delta_{T}}{2-\delta_{T}}\|f^{*}\|_{2}^{2}.$
For $\delta_{T}=\frac{1}{m^{2}}\mathcal{L}_{\mathrm{TE}}$, i.e. for stopping time $T=2\log m+\log\frac{1}{\mathcal{L}_{\mathrm{TE}}}$
${\mathcal{L}_{\mathrm{ST}}}\leq 8\alpha m(\log(\frac{1}{\delta}))^{2}\mathcal{L}_{\mathrm{TE}}\left(1+\frac{2\log^{2}m+\log^{2}\frac{1}{\mathcal{L}_{\mathrm{TE}}}}{m^{2}\cdot\alpha m(\left(\log\frac{1}{\delta}\right)^{2}}\right).$
As long as $\alpha m\left(\log\frac{1}{\delta}\right)^{2}=\Omega(m^{-2+\epsilon_{s}})$ for some $\epsilon>0$ this can be simplified to
${\mathcal{L}_{\mathrm{ST}}}\leq 8\alpha m(\log(\frac{1}{\delta}))^{2}\mathcal{L}_{\mathrm{TE}}\left(1+O(m^{-\epsilon_{s}})\right).$
Note that we can simplify the lower bound for $\mathcal{L}_{\mathrm{TE}}$, i.e. we have that with probabilit at least $1-2\exp(-\frac{d\delta^{2}}{4})-2\exp(-\frac{m\delta_{m}^{2}}{4})$
$\mathcal{L}_{\mathrm{TE}}\geq\frac{\alpha}{1+\alpha+\delta_{m}-\delta\alpha}-2\delta_{m}-\delta\alpha.$
As long as $\frac{1}{100}\geq\delta_{m}-\delta\alpha$, this can be further simplified to
$\mathcal{L}_{\mathrm{TE}}\geq 0.99\frac{\alpha}{1+\alpha}-2\delta_{m}-\delta\alpha.$
Therefore, taking $\delta=\frac{1}{d^{\frac{1}{2}-\epsilon_{d}}}$ and $\delta_{m}=\frac{1}{m^{\frac{1}{2}-\epsilon_{m}}}$, we have that with probability at least $1-2\exp(-\frac{1}{4}(\alpha m^{2})^{\epsilon_{d}})-2\exp(-\frac{1}{4}m^{\epsilon_{m}})$ that
$\mathcal{L}_{\mathrm{TE}}\geq 0.99\frac{\alpha}{1+\alpha}-2m^{-\frac{1}{2}+\epsilon_{m}}-(\alpha m^{2})^{-\frac{1}{2}+\epsilon_{d}}\alpha.$
Appendix E Weak-to-Strong Bound Applied to the Case of $2$-layer NN
This section deals with 2.2 and 6.4.
Proof of Theorem 3.1.
In Theorem E.2, if we select $\delta_{T}=\frac{1}{\sqrt{M_{\mathrm{TE}}}}\mathcal{L}_{\mathrm{TE}}$, then with stopping time $T=\frac{1}{\sigma_{k}^{2}}(\log M_{\mathrm{TE}}+\log\frac{1}{\mathcal{L}_{\mathrm{TE}}})$ we have that with probability $1-\frac{2}{M_{\mathrm{TE}}}-K\exp(-c\frac{M_{\mathrm{TE}}}{K})$
${\mathcal{L}_{\mathrm{ST}}}$ $\leq O_{d,k}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\mathcal{L}_{\mathrm{TE}}$
$+O_{d,k}\left(\frac{1}{\sigma_{k}^{4}}\frac{1}{\sqrt{M_{\mathrm{TE}}}}\right)\left(\log M_{\mathrm{TE}}-\log\mathcal{L}_{\mathrm{TE}}\right)^{2}\mathcal{L}_{\mathrm{TE}}+\frac{1}{\sqrt{M_{\mathrm{TE}}}}\mathcal{L}_{\mathrm{TE}}\|f^{*}\|_{\mathcal{D}}^{2}.$
This will hold if $M_{\mathrm{TE}}/\log M_{\mathrm{TE}}>K=N_{k}+\dots+N_{0}$ so it suffices to have $M_{\mathrm{TE}}\geq(N_{0}+\dots+N_{k})^{2}$. For our setup, this is $(N_{0}+\dots+N_{k})^{2}=\Theta(d^{2k})$. From this we have that
${\mathcal{L}_{\mathrm{ST}}}$ $\leq O_{d,k}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\mathcal{L}_{\mathrm{TE}}$
$+O_{d,k}\left(\frac{1}{\sigma_{k}^{4}}\frac{1}{\sqrt{M_{\mathrm{TE}}}}\right)\left(\log^{2}M_{\mathrm{TE}}+\log^{2}\mathcal{L}_{\mathrm{TE}}\right)\mathcal{L}_{\mathrm{TE}}+\frac{1}{\sqrt{M_{\mathrm{TE}}}}\mathcal{L}_{\mathrm{TE}}$
where we used the fact that the target is nomalized. Note now that from Lemma E.8, we have that $\sigma_{k}$ can be absorbed into the $O_{d,k}$. So we have that
${\mathcal{L}_{\mathrm{ST}}}$ $\leq O_{d,k}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\mathcal{L}_{\mathrm{TE}}$
$+O_{d,k}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\right)\log^{2}\mathcal{L}_{\mathrm{TE}}\mathcal{L}_{\mathrm{TE}}.$
This can be written as
${\mathcal{L}_{\mathrm{ST}}}\leq O_{d,k}\left(\frac{\log^{2}M_{\mathrm{TE}}+\log^{2}\mathcal{L}_{\mathrm{TE}}}{\sqrt{M_{\mathrm{TE}}}}\,\mathcal{L}_{\mathrm{TE}}\right).$
Now, to get the right exponent, we need a sharp lower bound on $\mathcal{L}_{\mathrm{TE}}$. We can get that under the Gaussian Universality Ansatz. Under Gaussian Universality Ansatz, the teacher’s test risk $\mathcal{L}_{\mathrm{TE}}$ behaves like the deterministic equivalent Equation 21. Therefore, Theorem D.6 applies, which tells us that
$\mathcal{L}_{\mathrm{TE}}\geq\Theta\left(M_{\mathrm{TE}}^{-\alpha}d^{2K}\|f\|_{\mathcal{D}}^{2}\right).$
In other words, $M_{\mathrm{TE}}^{-1}\leq O_{d,k}(\mathcal{L}_{\mathrm{TE}}^{\frac{1}{\alpha}})$.
Note that under Gaussian Universality, since $\mathcal{L}_{\mathrm{TE}}$ is lower bounded polyonomially in $M_{\mathrm{TE}}$, we can absorb $\log^{2}\mathcal{L}_{\mathrm{TE}}$ into $\log^{2}M_{\mathrm{TE}}$. Therefore, we get
${\mathcal{L}_{\mathrm{ST}}}$ $\leq O_{d,k}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\mathcal{L}_{\mathrm{TE}}.$
So we have that
${\mathcal{L}_{\mathrm{ST}}}\leq O_{d,k}\left(\frac{\log^{2}M_{\mathrm{TE}}}{\sqrt{M_{\mathrm{TE}}}}\right)\mathcal{L}_{\mathrm{TE}}\leq\tilde{O}_{d,k}\left(\mathcal{L}_{\mathrm{TE}}^{1+\frac{1}{2\alpha}}\right).$
Since we can take $\alpha$ to be close to $\frac{d+2}{d-2}$, for $d>200$ we will have this be $1.02$ so the final exponent is $1.49$. Finally, note that here $K$ is the number of nonzero $\sigma_{k}$ up to index $k$, so by Corollary E.9, it is of order $\Theta(d^{k})$. since $c$ is an absolute constant, for $M_{\mathrm{TE}}>\Theta(d^{2k})$, $K\exp(-c\frac{M_{\mathrm{TE}}}{d^{k}})<\frac{1}{M_{\mathrm{TE}}}$ and otherwise we can have the constant in $O_{d,k}$. ∎
Proof of Theorem 6.5.
This is given by Theorem E.1 for $t_{A}=\frac{1}{2}\sqrt{\frac{M_{\mathrm{TE}}}{K}}$. Note that since $M_{\mathrm{TE}}\geq K$, Assumption 6.2 holds with probability one. ∎
Theorem E.1 (Master Bound for Weak-to-Strong Generalization with 2-layer NN).
Under 2.2, if Assumption 6.1 is satisfied for $f^{*}$ and $K$. If $\sqrt{m}\geq t_{A}\sqrt{K}$ and the teacher attains minimum loss $\mathcal{L}_{\mathrm{TE}}=\hat{\mathcal{L}}_{\min}$, then the student trained to early stopping time $T=\frac{1}{\sigma_{K}^{2}}\log\frac{1}{\delta_{T}}$ will have with probability at least $1-2\delta-2K\exp(-ct_{A}^{2})$
${\mathcal{L}_{\mathrm{ST}}}$ $\leq\inf_{S\geq K}\Biggl{\{}\frac{m(\sum_{k=1}^{S}\sigma_{k}^{2})(\sum_{k=S+1}^{\infty}\sigma_{k}^{2})\log^{2}m}{\sigma_{S}^{4}(\sqrt{m}-t_{A}\sqrt{K})^{2}}\,\mathcal{L}_{\mathrm{TE}}+\frac{\sigma_{S+1}^{4}}{\sigma_{k}^{4}}\left(\log\frac{1}{\delta_{T}}\right)^{2}\,\mathcal{L}_{\mathrm{TE}}\Biggr{\}}+\frac{\delta_{T}}{2-\delta_{T}}\,\bigl{\|}f^{*}\bigr{\|}_{\mathcal{D}}^{2}.$
Proof of Theorem E.1.
Follows directly by instantiating Theorem 6.3 with the lower bound on $\kappa_{S}$ given by Theorem E.4 and the lower bound on $\lambda_{J}({\bm{A}})$ given by Theorem C.13. Finally, we just choose $T=\frac{1}{\sigma_{K}^{2}}\log\frac{1}{\delta_{T}}$. ∎
E.1 Weak-to-Strong Bound Applied to the Case of $2$-layer ReLU NN
Proof of Theorem 6.7.
This is given by Theorem E.2. ∎
Theorem E.2 (Weak-to-Strong Generalization with 2-layer ReLU NN).
Under 2.2, if Assumption 6.1 is satisfied for $f^{*}$ and $K$, let $k$ be the unique number such that $N_{0}+\dots+N_{k-1}+1\leq K\leq N_{0}+\dots+N_{k}$. If $M_{\mathrm{TE}}/\log M_{\mathrm{TE}}\geq K$ and the teacher attains minimum loss $\mathcal{L}_{\mathrm{TE}}=\hat{\mathcal{L}}_{\min}$, then the student trained to early stopping time $T=\frac{1}{\sigma_{k}^{2}}\log\frac{1}{\delta_{T}}$ will have with probability at least $1-\frac{2}{M_{\mathrm{TE}}}-K\exp(-c\frac{M_{\mathrm{TE}}}{k})$
${\mathcal{L}_{\mathrm{ST}}}$ $\leq O_{d,k}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\mathcal{L}_{\mathrm{TE}}$
$+O_{d,k}\left(\frac{1}{\sigma_{k}^{4}}\frac{1}{\sqrt{M_{\mathrm{TE}}}}\right)\left(\log\frac{1}{\delta_{T}}\right)^{2}\mathcal{L}_{\mathrm{TE}}+\frac{\delta_{T}}{2-\delta_{T}}\|f^{*}\|_{\mathcal{D}}^{2}.$
Here $c$ is an absolute constant.
Proof of Theorem E.2.
We will apply Theorem E.1. Note that the eigenvalues of the activation functions $\sigma$ in this case are $\sigma_{k}$ and note that for odd $k>1$, we have $\sigma_{k}=0$. Note that we can just ignore the zero eigenvalues. So in this case, $\sigma_{S}$ will actually be equal to $\sigma_{s}$ and $\sigma_{S+1}$ to $\sigma_{s+2}$ if we select $S=N_{0}+\dots+N_{s}$. Take $s$ from Proposition E.3. Then the proposition says that we have
$\frac{m(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k})(\sum_{k=s}^{\infty}\sigma_{k}^{2}N_{k})\left(\log\frac{1}{\delta}\right)^{2}}{\sigma_{s}^{4}\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}$ $\leq\Theta_{d}\left(\frac{m\sqrt{m}\left(\log\frac{1}{\delta}\right)^{2}}{\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}\right)$
$\lambda_{S+1}^{2}T^{2}=\frac{\sigma_{s+2}^{4}}{\sigma_{k}^{4}}\left(\log\frac{1}{\delta_{T}}\right)^{2}$ $\leq\Theta_{d}\left(\frac{1}{\sqrt{m}}\frac{1}{\sigma_{k}^{4}}\left(\log\frac{1}{\delta_{T}}\right)^{2}\right).$
Therefore, from Theorem E.1, we have that with probability at least $1-2\delta-2K\exp(-ct_{A}^{2})$
${\mathcal{L}_{\mathrm{ST}}}$ $\leq\Theta_{d}\left(\frac{m\sqrt{m}\left(\log\frac{1}{\delta}\right)^{2}}{\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}\right)\mathcal{L}_{\mathrm{TE}}$
$+\Theta\left(\frac{1}{\sqrt{m}}\frac{1}{\sigma_{k}^{2}}\left(\log\frac{1}{\delta_{T}}\right)^{2}\right)\left(\log\frac{1}{\delta_{T}}\right)^{2}\mathcal{L}_{\mathrm{TE}}+\frac{\delta_{T}}{2-\delta_{T}}\|f^{*}\|_{\mathcal{D}}^{2}.$
Take $\delta=\frac{1}{m}$ and take $t_{a}=m^{\frac{1}{2}}\frac{1}{2\sqrt{K}}$ so that $\sqrt{m}-t_{A}\sqrt{K}\geq\sqrt{m}/2$. Then $2K\exp(-ct_{A}^{2})\leq\frac{1}{m}$ if $\frac{m}{\log m}\geq K$. Note that therefore with probability at least $1-\frac{4}{m}$, we have
${\mathcal{L}_{\mathrm{ST}}}$ $\leq\Theta_{d}\left(\frac{1}{\sqrt{M_{\mathrm{TE}}}}\left(\log M_{\mathrm{TE}}\right)^{2}\right)\mathcal{L}_{\mathrm{TE}}$
$+\Theta\left(\frac{1}{\sigma_{k}^{4}}\frac{1}{\sqrt{M_{\mathrm{TE}}}}\right)\left(\log\frac{1}{\delta_{T}}\right)^{2}\mathcal{L}_{\mathrm{TE}}+\frac{\delta_{T}}{2-\delta_{T}}\|f^{*}\|_{\mathcal{D}}^{2}.$
E.2 Proofs Related to $2$-layer NN
Proposition E.3 (Optimal choice of $s$).
We can take $s=m^{\frac{1}{4(d-4)}}+o_{m}(1)$ ($s$ is closest even integer to $m^{\frac{1}{4(d-4)}}$) so that
$\frac{m(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k})(\sum_{k=s}^{\infty}\sigma_{k}^{2}N_{k})\left(\log\frac{1}{\delta}\right)^{2}}{\sigma_{s}^{4}\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}$ $\leq\Theta_{d}\left(\frac{m\sqrt{m}\left(\log\frac{1}{\delta}\right)^{2}}{\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}\right)$
$\lambda_{S+1}^{2}T^{2}=\frac{\sigma_{s+2}^{4}}{\sigma_{k}^{4}}\left(\log\frac{1}{\delta_{T}}\right)^{2}$ $\leq\Theta_{d}\left(\frac{1}{\sqrt{m}}\frac{1}{\sigma_{k}^{4}}\left(\log\frac{1}{\delta_{T}}\right)^{2}\right).$
Note that this $\Theta$ hides absolute constants and $d$ dependence.
Proof.
Note that for $k\gg d$ we have that $N_{k}=\Theta(k^{d-2})$. This follows directly from the closed form $N_{k}=\frac{(2k+d-2)(k+d-3)!}{k!(d-2)!}$. Furthermore, by Lemma E.8, we have that $\sigma_{k}^{2}=\Theta(k^{-(d+2)})$ for even $k\gg d$. Furthermore, $\sum_{k=1}^{\infty}\sigma_{k}^{2}N_{k}=\sigma(1)$. So, we have that
$\frac{m(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k})(\sum_{k=s}^{\infty}\sigma_{k}^{2}N_{k})\left(\log\frac{1}{\delta}\right)^{2}}{\sigma_{s}^{4}\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}$ $\leq\frac{1}{\sigma_{s}^{4}}\frac{m\sigma(1)^{2}\left(\log\frac{1}{\delta}\right)^{2}}{\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}.$
Therefore, taking $s=m^{\frac{1}{4(d+2)}}+o_{m}(1)$, where $o_{m}(1)$ is taken so that $s$ is the closest integer to $m^{\frac{1}{4(d+2)}}$. Then since $\sigma_{s}^{2}=\Theta(s^{-(d+2)})$, we have that $\sigma_{s}^{2}=\Theta(m^{-\frac{1}{4}})$, so $\frac{1}{\sigma_{s}^{4}}=\Theta(\sqrt{m})$. Note that $\sigma_{s+2}^{2}=\Theta((m^{\frac{1}{4(d+2)}}+2)^{-(d+2)})=\Theta(m^{-\frac{1}{4}})$. Therefore $\sigma_{s+2}^{4}=\Theta(\frac{1}{\sqrt{m}})$. ∎
Theorem E.4.
Under 2.2, if Assumption 6.1 is satisfied for $f^{*}$ and $K$, let $s$ be such that $N_{0}+\dots+N_{s-1}+1\leq K\leq N_{0}+\dots+N_{s}$. Then with probability at least $1-2\delta-2K\exp(-ct_{A}^{2})$
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq\frac{\|F_{B}^{\dagger}F_{A}\|_{\mathcal{H}}^{2}}{\lambda_{S}({\bm{A}})^{2}}\leq\frac{m(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k})(\sum_{k=s}^{\infty}\sigma_{k}^{2}N_{k})\left(\log\frac{1}{\delta}\right)^{2}}{\sigma_{s}^{4}\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}$
where $c$ is an absolute constant and $\delta\in(0,1)$. For $s\geq 2$ then withe probability at least $1-2\delta-2K\exp(-ct_{A}^{2})$
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq\frac{m\sigma(1)^{2}\left(\log\frac{1}{\delta}\right)^{2}}{\sigma_{s}^{4}\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}.$
Proof.
The first bound follows from Theorem C.13 applied to $\lambda_{J}({\bm{A}})$ and Proposition C.12 applied with Corollary E.7. This gives that with probability at least $1-2\delta-2K\exp(-ct_{A}^{2})$
$\lambda_{1}\!\left((\sqrt{{\bm{A}}_{S}})^{+}{\bm{B}}_{S}(\sqrt{{\bm{A}}_{S}})^{+}\right)\leq\frac{\|F_{B}^{\dagger}F_{A}\|_{\mathcal{H}}^{2}}{\lambda_{S}({\bm{A}})^{2}}\leq\frac{m(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k})(\sum_{k=s}^{\infty}\sigma_{k}^{2}N_{k})\left(\log\frac{1}{\delta}\right)^{2}}{\sigma_{s}^{4}\left(\sqrt{m}-t_{A}\sqrt{K}\right)^{4}}.$
The second bound follows by noting that $\sum_{k=1}^{\infty}\sigma_{k}^{2}N_{k}=\sigma(\langle{\bm{x}},{\bm{x}}\rangle)=\sigma(1)$. ∎
In this case, the basis of eigenfunctions can be taken to be the spherical harmonics $\{\{\phi_{k,i}\}_{i=1}^{N_{k}}\}_{k=1}^{\infty}$ where $\phi_{k,i}$ is the $i$-th spherical harmonic of order $k$ in $d$ dimensions and $N_{k}=\frac{(2k+d-2)(k+d-3)!}{k!(d-2)!}$ is the dimensionality of spherical harmonics of order $k$ in $d$ dimensions.
Proposition E.5 (Eigendecomposition of the kernel induced by random features in the case of $2$-layer NN).
Under 2.2, the kernel induced by the random features is given by
$\mathcal{K}({\bm{x}},{\bm{x}}^{\prime})=\sum_{k=1}^{\infty}\sigma_{k}^{2}\langle\phi_{k}({\bm{x}}),\phi_{k}({\bm{x}}^{\prime})\rangle.$
Here, $\phi_{k}:\mathbb{R}^{d}\to\mathbb{R}^{N_{k}}$ is a vector of spherical harmonics of order $k$ in $d$ dimension, $N_{k}=\frac{(2k+d-2)(k+d-3)!}{k!(d-2)!}$, and $\sigma_{k}$ is the $k$-th term in the decomposition of the ReLU functions $\sigma$ to the basis of $\{\phi_{k}\}_{i=1}^{\infty}$.
We provide the estimates for the size of $\sigma_{k}$ in Section E.3.
Finally, we find $B$ and $L$ in the spherical case.
Proposition E.6.
Under 2.2, if $J=N_{0}+N_{1}+\dots+N_{s}$ then
$\|P_{\leq J}g\|_{\mathcal{D}}^{2}$ $=\sum_{k=1}^{s}\sigma_{k}^{2}N_{k}$
$\|P_{\geq J+1}g\|_{\mathcal{D}}^{2}$ $=\sum_{k=s+1}^{\infty}\sigma_{k}^{2}N_{k}.$
So $L=B=\sqrt{\left(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k}\right)\left(\sum_{k=s+1}^{\infty}\sigma_{k}^{2}N_{k}\right)}$.
Proof.
We have that
$P_{\leq J}g$ $=\sum_{k=1}^{s}\sigma_{k}\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{x}})$
$\|P_{\leq J}g\|_{\mathcal{D}}^{2}$ $=\langle\sum_{k=1}^{s}\sigma_{k}\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{x}}),\sum_{l=1}^{s}\sigma_{l}\phi_{l}({\bm{w}})^{T}\phi_{l}({\bm{x}})\rangle_{\mathcal{D}}=\sum_{k=1}^{s}\sum_{l=1}^{s}\sigma_{k}\sigma_{l}\langle\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{x}}),\phi_{l}({\bm{w}})^{T}\phi_{l}({\bm{x}})\rangle_{\mathcal{D}}.$
Now since $\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{x}})=\sum_{j=1}^{N_{k}}\phi_{k,j}({\bm{w}})\phi_{k,j}({\bm{x}})$ and since $\{\phi_{k,j}\}$ are orthogonal w.r.t $\mathcal{D}$, only the terms of the form $\phi_{k,j}({\bm{w}})\phi_{k,j}({\bm{x}})\phi_{k,j}({\bm{w}})\phi_{k,j}({\bm{x}})$ remain. For these, we have that $\mathbb{E}_{\mathcal{D}}\left(\phi_{k,j}({\bm{w}})\phi_{k,j}({\bm{x}})\phi_{k,j}({\bm{w}})\phi_{k,j}({\bm{x}})\right)=\phi_{k,j}({\bm{w}})^{2}\cdot 1$. So we have that
$\|P_{\leq J}g\|_{\mathcal{D}}^{2}$ $=\sum_{k=1}^{s}\sigma_{k}^{2}\langle\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{x}}),\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{x}})\rangle_{\mathcal{D}}$
$=\sum_{k=1}^{s}\sigma_{k}^{2}\phi_{k,j}({\bm{w}})^{2}=\sum_{k=1}^{s}\sigma_{k}^{2}N_{k}.$
The last step holds because $\phi_{k}({\bm{w}})^{T}\phi_{k}({\bm{w}})=N_{k}P_{k}(\langle{\bm{w}},{\bm{w}}\rangle)=N_{k}P_{k}(1)=N_{k}$.
The same argument shows that $\|P_{J+1}g\|_{\mathcal{D}}^{2}=\sum_{k=s+1}^{\infty}\sigma_{k}^{2}N_{k}$. ∎
Corollary E.7.
Under 2.2, if $N_{0}+\dots+N_{s-1}+1\leq J\leq N_{0}+\dots+N_{s}$ then
$L=B=\sqrt{(\sum_{k=1}^{s}\sigma_{k}^{2}N_{k})(\sum_{k=s}^{\infty}\sigma_{k}^{2}N_{k})}.$
This simplifies to $L=B\leq\Theta(\frac{1}{d})\sqrt{\frac{2^{s}}{s!}}$.
Proof.
Note that if $J=N_{0}+N_{1}+\dots+N_{s-1}+p$ for some $1\leq p\leq N_{s}$ then we have that
$\|P_{\leq J}g\|_{\mathcal{D}}^{2}$ $=\sum_{k=1}^{s-1}\sigma_{k}^{2}N_{k}+p\sigma_{s}^{2}$
$\|P_{\geq J+1}g\|_{\mathcal{D}}^{2}$ $=(N_{k}-p)\sigma_{s}^{2}+\sum_{k=s+1}^{\infty}\sigma_{k}^{2}N_{k}.$
E.3 Computation of the Kernel Eigenvalues for the case $2$-layer NN
Proof.
of Proposition E.5: For ReLU activation, we can do the following decomposition:
$\sigma(\langle{{\bm{w}}},{{\bm{x}}}\rangle)=\sum_{k\geq 0}N_{k}\sigma_{k}P_{d,k}(\langle{{\bm{w}}},{{\bm{x}}}\rangle)=\sum_{k\geq 0}\sigma_{k}\langle{{\bm{\phi}}_{d,k}({\bm{w}})},{{\bm{\phi}}_{d,k}({\bm{x}})}\rangle,$
Here $\sigma_{k}$ is given by
$\sigma_{k}=\frac{\lvert\mathbb{S}^{d-2}\rvert}{\lvert\mathbb{S}^{d-1}\rvert}\int_{0}^{1}sP_{k,d}(s)(1-s^{2})^{\frac{d-3}{2}}\mathrm{d}s.$
Combining this with $\lvert\mathbb{S}^{d-1}\rvert=\frac{2\pi^{d/2}}{\Gamma(\frac{d}{2})}$, we have
$\sigma_{k}=\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\int_{0}^{1}sP_{k,d}(s)(1-s^{2})^{\frac{d-3}{2}}\mathrm{d}s.$
The eigendecomposition of the kernel in this case is given by
$\mathcal{K}({\bm{x}},{\bm{x}}^{\prime})=\mathbb{E}_{{\bm{w}}}[\sigma(\langle{{\bm{w}}},{{\bm{x}}}\rangle)\sigma(\langle{{\bm{w}}},{{\bm{x}}^{\prime}}\rangle)]=\mathbb{E}_{{\bm{w}}}\left[\sum_{k\geq 0}\sigma_{k}\langle{{\bm{\phi}}_{k}({\bm{w}})},{{\bm{\phi}}_{k}({\bm{x}})}\rangle\cdot\sigma_{k}\langle{{\bm{\phi}}_{k}({\bm{w}})},{{\bm{\phi}}_{k}({\bm{x}}^{\prime})}\rangle\right]=\sum_{k\geq 0}\sigma_{k}^{2}\langle{{\bm{\phi}}_{k}({\bm{x}})},{{\bm{\phi}}_{k}({\bm{x}}^{\prime})}\rangle.$
Now, we find the size of $\sigma_{k}$.
Lemma E.8 (Size of $\sigma_{k}$).
Depending on the size of $d$ and $k$, the following holds for the size of coefficients of ReLU activation in the basis of Legendre polynomials $\sigma_{k}=\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\cdot\frac{1}{2^{k}}\frac{1}{(\frac{k}{2}+\frac{d-1}{2})\dots(\frac{d-1}{2})}$ for even $k$, $\sigma_{1}=\frac{1}{2d}$, and $\sigma_{k}=0$ for odd $k>1$. Asymptotically, this means
  1. 1.
    $\sigma_{0}=\frac{1}{\sqrt{2\pi d}}+\Theta(d^{-\frac{3}{2}})$, $\sigma_{1}=\frac{1}{2d}$
  2. 2.
    If $k=\Theta(1)$, then for even $k$, $\sigma_{k}=\Theta(d^{-\frac{k+1}{2}})$.
  3. 3.
    If $k\gg d$, then for even $k$, $\sigma_{k}=\Theta(k^{-\frac{d-1}{2}-\frac{3}{2}})$.
Proof.
The third case is shown in Petrini et al. (2022). To prove the general claim, we use the formula derived in Petrini et al. (2022)
$\sigma_{k}=\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\cdot\frac{\Gamma(\frac{d-1}{2})}{\Gamma(k+\frac{d-1}{2})}\left(-\left(-\frac{1}{2}\right)^{k}\right)\frac{d^{k-2}}{dt^{k-2}}(1-t^{2})^{k+\frac{d-3}{2}}\Big{|}_{0}^{1}.$
Note that all terms will be multiplied by $(1-t^{2})^{2+\frac{d-3}{2}}$ which is 00 at $t=1$, so we only need to consider $t=0$. Note that if a term has a nonzero power of $t$, it will be evaluated to 00. When taking $j$-th ($j\leq k-2$) derivative of the expression above, all the terms will be in the form of $(1-t^{2})^{l+\frac{d-3}{2}}t^{s}$ times some constant that depends on $d$ for some integers $l$ and $s$. So if out of $k-2$ derivatives we take $a$ on the term with $(1-t^{2})$ and $k-2-a$ on $t^{s}$ (note that the ordering doesn’t matter as we only add constants), in order to have term with no $t$ at the end, we need to have $k-2-a=s$, but $s=a$ so we have that $a=\frac{k-2}{2}$. Therefore, the evaluated derivate equals exactly
$\frac{d^{k-2}}{dt^{k-2}}(1-t^{2})^{k+\frac{d-3}{2}}\Big{|}_{0}^{1}=\left(k+\frac{d-3}{2}\right)\dots\left(k-(\frac{k-2}{2}-1)+\frac{d-3}{2}\right)=\left(k+\frac{d-3}{2}\right)\dots\left(\frac{k+4}{2}+\frac{d-3}{2}\right).$
Note that $\frac{\Gamma(\frac{d-1}{2})}{\Gamma(k+\frac{d-1}{2})}=\frac{1}{(k-1+\frac{d-1}{2})\dots(\frac{d-1}{2})}$, therefore we can compute that
$\sigma_{k}$ $=\frac{1}{2^{k}}\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\cdot\frac{1}{(k-1+\frac{d-1}{2})\dots(\frac{d-1}{2})}\left(k+\frac{d-3}{2}\right)\dots\left(\frac{k+4}{2}+\frac{d-3}{2}\right)$
$\sigma_{k}$ $=\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\cdot\frac{1}{2^{k}}\frac{1}{(\frac{k}{2}+\frac{d-1}{2})\dots(\frac{d-1}{2})}.$
$\sigma_{k}$ $=\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\cdot\frac{1}{2}\frac{1}{(k+d-1)((k-2)+d-1)((k-4)+d-1)\dots(2+d-1)(d-1)}.$
So, when $k=\Theta(1)$ we have that this is $\sigma_{k}=\Theta(d^{\frac{1}{2}}d^{-\frac{k+2}{2}})=\Theta(d^{-\frac{k+1}{2}})$, since the first term is $\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}=\Theta(d^{\frac{1}{2}})$. ∎
Corollary E.9 (Size of $\sigma_{k}^{2}N_{k}$).
The following bound holds for $\sigma_{k}^{2}N_{k}$ for $k$ even or $k=1$ If $k\gg d$ then $\sigma_{k}^{2}N_{k}=\Theta(k^{-4})$. If $k\ll d$ then $\sigma_{k}^{2}N_{k}=\Theta(d^{-1})$. If $k>1$ is odd, then $\sigma_{k}^{2}N_{k}=0$.
Proof.
We have for $\sigma_{k}^{2}N_{k}$ that
$\sigma_{k}^{2}N_{k}$ $=\frac{1}{4}\left(\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\right)^{2}\cdot\frac{(d-1)(d-1+1)\dots(d-1+k-2)}{(k+d-1)^{2}\dots(d-1+2)^{2}(d-1)^{2}}(2k+d-2)$
$=\frac{1}{4}\left(\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\right)^{2}\cdot(2k+d-2)\frac{(d-1+1)(d-1+3)\dots(d-1+k-3)}{(d-1)(d-1+2)\dots(d-1+k-4)(d-1+k-2)}\frac{1}{(d-1+k)^{2}}.$
When $k\ll d$, we can just count the terms with $d$ in the numerator and denominator and note that $\left(\frac{\Gamma(\frac{d}{2})}{\sqrt{\pi}\Gamma(\frac{d-1}{2})}\right)^{2}=\Theta(d)$. For $k\gg d$, we just need to count the terms with $k$ in the numerator and denominator. The results follow.
Appendix F Proof of Weak-to-Strong Limitation
Proof of Theorem 4.1.
This follows from Theorem 6.11 by noting that from Lemma 6.10 the teacher satisfies Definition 6.9 with respect to the target and the student satisfies Definition 6.9 with respect to the teacher. ∎
Proof of Lemma 6.10.
From Equation 13, the gradient flow solution is $f_{T}=\sum_{k\geq 1}(1-e^{-\lambda_{k}T})\langle f^{*},e_{k}\rangle_{\mathcal{D}}e_{k}$. For shrinkage optimality, we need to show that $\alpha=1$ minimizes $\mathbb{E}_{\mathbf{x}}[(\alpha f_{T}(\mathbf{x})-f^{*}(\mathbf{x}))^{2}]$ for $0\leq\alpha\leq 1$.
Expanding this expression in the eigenbasis:
$\mathbb{E}_{\mathbf{x}}[(\alpha f_{T}(\mathbf{x})-f^{*}(\mathbf{x}))^{2}]$ $=\left\|\alpha\sum_{k\geq 1}(1-e^{-\lambda_{k}T})\langle f^{*},e_{k}\rangle_{\mathcal{D}}e_{k}-\sum_{k\geq 1}\langle f^{*},e_{k}\rangle_{\mathcal{D}}e_{k}\right\|^{2}_{\mathcal{D}}$ (19)
$=\sum_{k\geq 1}(\alpha(1-e^{-\lambda_{k}T})-1)^{2}\langle f^{*},e_{k}\rangle^{2}_{\mathcal{D}}$ (20)
The proof is immediate by noting that $1-e^{-\lambda_{k}T}\geq 0$ and $\alpha(1-e^{-\lambda_{k}T})\leq 1$ for all $k$ and $0\leq\alpha\leq 1$. ∎
Proof of Theorem 6.11.
The proof is a straightforward geometric argument in the function space. Note first that Definition 6.9 gives a restiction of where the teacher predictor $f_{\text{teacher}}$ can be in the function space with respect to the target $f^{*}$ and the origin. Note that $\|f_{\text{teacher}}-f^{*}\|_{\mathcal{D}}^{2}\leq\|\alpha f_{\text{teacher}}-f^{*}\|_{\mathcal{D}}^{2}$ for all $0\leq\alpha\leq 1$ holds if and only if $\langle{f_{\text{teacher}}},{f_{\text{teacher}}-f^{*}}\rangle\leq 0$. Similarlry, LABEL:ass:shrinkageopimalitystudent holds if and only if $\langle{f_{\text{student}}},{f_{\text{teacher}}-f_{\text{student}}}\rangle\leq 0$. This implies that the teacher predictor $f_{\text{teacher}}$ is inside the set $f_{\text{teacher}}\in\{f\mid\|f-\frac{f^{*}}{2}\|_{\mathcal{D}}\leq\frac{1}{2}\|f^{*}\|_{\mathcal{D}}\}$. Similarly, we have that $f_{\text{student}}\in\{f\mid\|f-\frac{f_{\text{teacher}}}{2}\|_{\mathcal{D}}\leq\frac{1}{2}\|f_{\text{teacher}}\|_{\mathcal{D}}\}$. Both of these sets are spheres in the function space w.r.t. $\mathcal{D}$-norm. Let $\beta\geq 1$ be such that $\tilde{f}_{\text{teacher}}=\beta f_{\text{teacher}}$ and such that $\langle{\tilde{f}_{\text{teacher}}},{f^{*}-\tilde{f}_{\text{teacher}}}\rangle=0$. Then we have that
$\mathcal{L}_{\mathrm{TE}}=1-\beta^{2}f_{\text{teacher}}^{2}+(\beta-1)^{2}f_{\text{teacher}}^{2}=1-(2\beta-1)f_{\text{teacher}}^{2}.$
Let $\tilde{f}_{\text{student}}$ the predictor inside the student set with the smallest risk, i.e. on the intersection of the line connecting $f^{*}$ and $\frac{f_{\text{teacher}}}{2}$ and the boundary of the allowed student set. Note then that
$\mathcal{L}(\tilde{f}_{\text{student}})=\left(\|f^{*}-\frac{f_{\text{teacher}}}{2}\|_{\mathcal{D}}-\|\frac{f_{\text{teacher}}}{2}\|_{\mathcal{D}}\right)^{2}$
Note further that
$\|f^{*}-\frac{f_{\text{teacher}}}{2}\|_{\mathcal{D}}^{2}=1-\beta^{2}f_{\text{teacher}}^{2}+(\beta-\frac{1}{2})^{2}f_{\text{teacher}}^{2}=1-(\beta-\frac{1}{4})f_{\text{teacher}}^{2}.$
Therefore, we have that
$f_{\text{teacher}}^{2}=\frac{1-\mathcal{L}_{\mathrm{TE}}}{2\beta-1}.$
From this we get that
$\mathcal{L}(\tilde{f}_{\text{student}})$ $=\left(\sqrt{1-(\beta-\frac{1}{4})\frac{1-\mathcal{L}_{\mathrm{TE}}}{2\beta-1}}-\frac{1}{2}\sqrt{\frac{1-\mathcal{L}_{\mathrm{TE}}}{2\beta-1}}\right)^{2}$
$=\frac{1}{8\beta-4}\left(\sqrt{4\beta-3+(4\beta-1)\mathcal{L}_{\mathrm{TE}}}-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}$
Note that for $\beta\geq 1$ and $\mathcal{L}_{\mathrm{TE}}\leq 1$ we have that
$\mathcal{L}(\tilde{f}_{\text{student}})$ $=\frac{1}{8\beta-4}\left(\sqrt{4\beta-3+(4\beta-1)\mathcal{L}_{\mathrm{TE}}}-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}$
$\geq\frac{1}{2}\left(1+\mathcal{L}_{\mathrm{TE}}-\sqrt{1+2\mathcal{L}_{\mathrm{TE}}-3\mathcal{L}_{\mathrm{TE}}^{2}}\right).$
The last inequality holds by noticiting that the function is increasing in $\beta$, so it is minimized for $\beta=1$ for $10<\mathcal{L}_{\mathrm{TE}}<$ 1.
That is, since ${\mathcal{L}_{\mathrm{ST}}}\geq\mathcal{L}(\tilde{f}_{\text{student}})$ by design, we have that
${\mathcal{L}_{\mathrm{ST}}}\geq\frac{1}{2}\left(1+\mathcal{L}_{\mathrm{TE}}-\sqrt{1+2\mathcal{L}_{\mathrm{TE}}-3\mathcal{L}_{\mathrm{TE}}^{2}}\right)\geq\frac{3}{4}\mathcal{L}_{\mathrm{TE}}^{2}.$
The last inequality $\frac{1}{2}\left(1+\mathcal{L}_{\mathrm{TE}}-\sqrt{1+2\mathcal{L}_{\mathrm{TE}}-3\mathcal{L}_{\mathrm{TE}}^{2}}\right)\geq\frac{3}{4}\mathcal{L}_{\mathrm{TE}}^{2}$ holds for $0\leq\mathcal{L}_{\mathrm{TE}}\leq 1$. ∎
Proof of Theorem 6.12.
The proof is similar to the proof of Theorem 6.11. Similarly, we have that the teacher predictor is bounded to the set $f_{\text{teacher}}\in\{f\mid\|f-\frac{f^{*}}{2}\|_{\mathcal{D}}\leq\frac{1}{2}\|f^{*}\|_{\mathcal{D}}\}$. The student predictor is in this case bounded to the set $f_{\text{student}}\in\{f\mid\|f\|_{\mathcal{D}}\leq\|f_{\text{teacher}}\|_{\mathcal{D}}\}$. Let $\beta=\langle{f^{*}-f_{\text{teacher}}},{f^{*}}\rangle$. Note that from the condition $f_{\text{teacher}}\in\{f\mid\|f-\frac{f^{*}}{2}\|_{\mathcal{D}}\leq\frac{1}{2}\|f^{*}\|_{\mathcal{D}}\}$ we have that $\langle{f_{\text{teacher}}},{f^{*}-f_{\text{teacher}}}\rangle\leq 0$ which implies that there is $\beta\geq 1$ for which we have $\langle{\beta f_{\text{teacher}}},{f^{*}-f_{\text{teacher}}}\rangle$. Note that $\mathcal{L}_{\mathrm{TE}}=\|f^{*}-f_{\text{teacher}}\|_{\mathcal{D}}^{2}$. Therefore, we have that since $\|f^{*}\|_{\mathcal{D}}^{2}=1$
$\mathcal{L}_{\mathrm{TE}}=1-\beta^{2}f_{\text{teacher}}^{2}+(\beta-1)^{2}f_{\text{teacher}}^{2}=1-(2\beta-1)f_{\text{teacher}}^{2}.$
This implies that
$f_{\text{teacher}}^{2}=\frac{1-\mathcal{L}_{\mathrm{TE}}}{2\beta-1}.$
Note that
${\mathcal{L}_{\mathrm{ST}}}\geq(\|f^{*}\|_{\mathcal{D}}-\|f_{\text{teacher}}\|_{\mathcal{D}})^{2}=(1-\|f_{\text{teacher}}\|_{\mathcal{D}})^{2}$
by the same argument as in Theorem 6.11. Therefore, we have that
${\mathcal{L}_{\mathrm{ST}}}\geq\left(1-\sqrt{\frac{1-\mathcal{L}_{\mathrm{TE}}}{2\beta-1}}\right)^{2}$
Note that $1-\sqrt{\frac{1-\mathcal{L}_{\mathrm{TE}}}{2\beta-1}}$ is positive and increasing in $\beta$, so the above lower bound is minimized for $\beta=1$. Therefore
${\mathcal{L}_{\mathrm{ST}}}\geq\left(1-\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\right)^{2}=2-\mathcal{L}_{\mathrm{TE}}-2\sqrt{1-\mathcal{L}_{\mathrm{TE}}}\geq\frac{1}{4}\mathcal{L}_{\mathrm{TE}}^{2}.$
Proof of Corollary 6.13.
The proof is immediate from Theorem 6.12 and the fact that the student $f_{\text{student}}^{(i)}$ is shrinking-optimal w.r.t. $f_{\text{student}}^{(i-1)}$ implies that $\|f_{\text{student}}^{(i)}\|_{\mathcal{D}}\leq\|f_{\text{student}}^{(i-1)}\|_{cD}$, which further implies that $\|f_{\text{student}}^{(i)}\|_{\mathcal{D}}\leq\|f_{\text{teacher}}\|_{\mathcal{D}}$. ∎
Appendix G Deterministic Equivalent of the Teacher Error
Let ~$\tilde{\mathcal{L}}(f^{*},m)$ be the deterministic equivalent of the test risk of our model with $m$ random features trained on population loss with target $f^{*}$. Let $s_{K}$ be the eigenvalues of the activation function, i.e. the decompositon of the activation function in the eigenbasis $\{e_{i}\}_{i=1}^{\infty}$. Let $\beta^{*}$ be the coefficients of the target $f^{*}$ in the eigenbasis, $f^{*}=\sum_{i=1}^{\infty}\beta_{i}^{*}e_{i}$. To write down a closed form of ~$\tilde{\mathcal{L}}(f^{*},m)$, we introduce the following notation ${\bm{S}}$ and $\nu$
${\bm{S}}$ $=\text{diag}(s_{1},s_{2},\dots)$
$m$ $=\text{tr}({\bm{S}}({\bm{S}}+\nu)^{-1}).$
That is, $\nu$ is the unique solution to this equation. According to Corollary $3.5$ in Defilippis et al. (2024), the deterministic equivalent of test risk has the following closed form then
~$\tilde{\mathcal{L}}(f^{*},m)=\nu\langle\beta^{*},({\bm{S}}+\nu)^{-1}\beta^{*}\rangle.$ (21)
G.1 Lower Bound for Teacher Error
We show a lower bound on the deterministic equivalent of test risk of a Random Feature Model with $m$ features that we consider in our setup. Under the Gaussian Universality Ansatz, the true test error of the teacher’s predictor $\mathcal{L}_{\mathrm{TE}}$ will behave like its deterministic equivalent.
Theorem G.1 (Lower bound for the Error of a $2$-layer random feature ReLU).
If $f^{*}$ satisfies Assumption 6.1 for 2.3 with $K$ such that for some $k$, $N_{0}+\dots+N_{k-1}+1\leq K\leq N_{0}+\dots+N_{k}$, then for fixed $d$ there exists $\alpha>\frac{d+2}{d-2}$ such that for the deterministic equivalent of the test risk we have
~$\tilde{\mathcal{L}}(f^{*},m)\geq\Theta\left(\frac{1}{m^{\alpha}}d^{2}\|f^{*}\|_{\mathcal{D}}^{2}\right).$
We can take $\alpha=\frac{d+2}{d-2}+o_{d}(1)$.
Proof of Theorem G.1.
Note that the eigenvalues of the activation function are $\sigma_{s}^{2}$ with multiplicity $N_{s}$. We want to lower bound $\nu$. Assume that for large $m$, $\nu\leq\frac{1}{m^{\alpha}}$. Note that we have the following two things for $s\gg d$,
$\sigma_{s}^{2}=\Theta(s^{-(d+2)})\text{ and }N_{s}=\Theta(s^{d-2}).$
The first follows from Lemma E.8 and the second one follows immediately from the closed form of $N_{s}$. Note that in this case we have that $s_{i}=\sigma_{i}^{2}$. Note that then for $\sigma_{s}^{2}>\nu$ we have that $\frac{\sigma_{s}^{2}}{\sigma_{s}^{2}+\nu}>\frac{1}{2}$. Note that if $\nu<\frac{1}{m^{\alpha}}$ then we have that for all $s<m^{\frac{\alpha}{d+2}}$ that $\sigma_{s}^{2}>\nu$. Note that there is at least $N_{s_{\max}}$ of those, where $s_{\max}$ is the largest such $s$. We have that $s_{\max}\geq m^{\frac{\alpha}{d+2}}$ But we have that $N_{s_{\max}}=\Theta(s_{\max}^{d-2})\geq\Theta(m^{\alpha\frac{d-2}{d+2}})$. Note also that with this choice of $s_{\max}$
$m=\text{tr}({\bm{S}}({\bm{S}}+\nu)^{-1})=\sum_{i=1}^{\infty}\frac{s_{i}}{s_{i}+\nu}=\sum_{s=1}^{\infty}\sum_{i=1}^{N_{s}}\frac{\sigma_{s,i}^{2}}{\sigma_{s,i}^{2}+\nu}\geq N_{s_{\max}}\frac{1}{2}\geq\Theta(m^{\alpha\frac{d-2}{d+2}}),$
where $\sigma_{s,i}=\sigma_{s}$.
So if $\alpha=\frac{d+2}{d-2}+o_{d}(1)$ then we have a contradiction. Therefore $\nu\geq\frac{1}{m^{\alpha}}$ for this $\alpha=\frac{d+2}{d-2}+o_{d}(1)$. Plugging this back in we have
~$\tilde{\mathcal{L}}(f^{*},m)=\nu\left(\sum_{i=1}^{\infty}\frac{1}{\sigma_{i}^{2}+\nu}\sum_{j=1}^{N_{i}}\beta_{i,j}^{*2}\right)\geq\frac{1}{2}\nu\frac{1}{\sigma_{1}^{2}+\frac{1}{m}}\|f^{*}\|_{\mathcal{D}}^{2}.$
For the last inequality, note first that $\sigma_{i}\leq\sigma_{1}$ so $\frac{1}{\sigma_{i}^{2}+\nu}\geq\frac{1}{\sigma_{1}^{2}+\nu}$. Also note that
$m=\sum_{i=1}^{\infty}\frac{s_{i}}{s_{i}+\nu}\leq\frac{\sum_{i=1}^{\infty}s_{i}}{\nu},$
so $\nu\leq\frac{\sum_{i=1}^{\infty}s_{i}}{m}$. Note that since $\sum_{s=1}^{\infty}\sigma_{s}^{2}N_{s}=\sigma(1)=1$ we have that $\sum_{i=1}^{\infty}s_{i}=1$. Therefoe, it holds that $\nu\leq\frac{1}{m}$. This finishes the proof. ∎
G.2 No $\Theta(1)$-error Asymptotics for ReLU Networks
Proof of Theorem 3.4.
This follows from Theorem G.2. Note that under Gaussian Universality, this theorem holds for the test risk. ∎
Theorem G.2 (Upper Bound on Teacher Error with ReLU Network).
If $f^{*}$ is spanned by the spherical harmonics of order that is even and at mos $k$ or $1$, for 2.2 in fixed dimension $d$, we have the following bound on the deterministic equivalent of the test risk
~$\tilde{\mathcal{L}}(f^{*},m)\leq\Theta\left(\frac{1}{m}d^{(k+1)}\|f^{*}\|_{\mathcal{D}}^{2}\right)$
Proof of Theorem G.2.
Note that
$m=\text{tr}({\bm{S}}({\bm{S}}+\nu)^{-1})=\sum_{i=1}^{\infty}\frac{s_{i}}{s_{i}+\nu}\leq\frac{\sum_{i=1}^{\infty}s_{i}}{\nu}$
so we have $\nu\leq\frac{\sum_{i=1}^{\infty}s_{i}}{m}$. Since $\sum_{i=1}^{\infty}s_{i}=\sigma(1)=1$, we have that $\nu\leq\frac{1}{m}$. Then we have that since $f^{*}$ is spanned by harmonics of order at most $k$
~$\tilde{\mathcal{L}}(f^{*},m)=\nu\left(\sum_{i=1}^{\infty}\frac{1}{\sigma_{i}^{2}+\nu}\sum_{j=1}^{N_{i}}\beta_{i,j}^{*2}\right)=\nu\sum_{i=1}^{k}\frac{1}{\sigma_{i}^{2}+\nu}\sum_{j=1}^{N_{s}}\beta_{i,j}^{*2}\leq\frac{\nu}{\sigma_{k}^{2}+\nu}\sum_{i=1}^{k}\sum_{j=1}^{N_{i}}\beta_{i}^{*2}=\frac{\nu}{\sigma_{k}^{2}+\nu}\|f^{*}\|_{\mathcal{D}}^{2}.$
Finally, note that $\sigma_{k}^{2}=\Theta(d^{-(k+1)})$ so we have $\frac{\nu}{\nu+\sigma_{k}^{2}}\leq\frac{1}{m}\Theta(d^{(k+1)})$
Remark G.3.
The same proof shows that in the Gaussian Features B.1 if the covairance structure is fixed we cannot have a $\Theta(1)$-error asymptotic for any function that is learnable by the model as $M_{\mathrm{TE}}\to\infty$,.
Appendix H Experiment Details
In this section, we provide experiment details for simulating weak-to-strong generalization for both 2-layer ReLU networks (Figure 2) and linear networks (Figure 3). All experiments are conducted on one H100 GPU, and are finished within 24 GPU hours.
To avoid the effect of randomness, experiments are repeated $5$ times with different random seeds, 95% confidence intervals are shown in Figure 2(a) and Figure 3(a).
For curve fitting in Figure 2(b) and Figure 3(b), we use numpy.polyfit with degree $1$ after taking the log of teacher loss $\mathcal{L}_{\mathrm{TE}}$ and student loss ${\mathcal{L}_{\mathrm{ST}}}$.
For linear network, we simulate the limit case where student size $M_{\mathrm{ST}}\to\infty$. The simulation follows Equation 13 with $\lambda_{1}=1$, $\lambda_{i}=(d-1)^{-2/3}$ for $i>1$. Note this method is not viable for 2-layer ReLU network because $\lambda_{i}$ is long-tailed for 2-layer ReLU network. Therefore, we set $M_{\mathrm{ST}}=16384$ in Figure 2.
Appendix I Additional Experimental Results
I.1 2-Layer ReLU Network Experiments
Figure 4 shows full experimental results for 2.2. It additionally includes $d=16$ and $64$ compared to Figure 2. We observe that when $d=64$, the student loss ${\mathcal{L}_{\mathrm{ST}}}$ is polynomially smaller than the teacher loss $\mathcal{L}_{\mathrm{TE}}$, with a higher estimated exponent compared to $d=32$. When $d=16$, the curve suggests a sudden change of optimal early stopping time, which we believe is caused by lack of student size $M_{\mathrm{ST}}$.
However, due to GPU memory limit, we are not able to increase $M_{\mathrm{ST}}$. Thus we conduct ablation study with smaller $M_{\mathrm{ST}}$ to investigate the impact of $M_{\mathrm{ST}}$. In Figure 5, we compare the loss ratio curve between $M_{\mathrm{ST}}=8192$ and $M_{\mathrm{ST}}=16384$. We observe larger gap between the two when teacher loss $\mathcal{L}_{\mathrm{TE}}$ is smaller, i.e., for small $d$ and large $M_{\mathrm{TE}}$. Therefore, it is not hard to believe that ${\mathcal{L}_{\mathrm{ST}}}=\tilde{\Theta}(\mathcal{L}_{\mathrm{TE}}^{2})$ for this setting as well. We leave it as future work.
(a) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}$ v.s. $t$ at $d=16$
(b) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}$ v.s. $t$ at $d=32$
(c) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}$ v.s. $t$ at $d=64$
(d) ${\mathcal{L}_{\mathrm{ST}}}$ v.s. $\mathcal{L}_{\mathrm{TE}}$ at $d=16$
(e) ${\mathcal{L}_{\mathrm{ST}}}$ v.s. $\mathcal{L}_{\mathrm{TE}}$ at $d=32$
(f) ${\mathcal{L}_{\mathrm{ST}}}$ v.s. $\mathcal{L}_{\mathrm{TE}}$ at $d=64$
Figure 4: Weak-to-strong generalization happens in 2-layer ReLU networks with input dimension $d=16,32,64$, student size $M_{\mathrm{ST}}=16384$, and teacher size $M_{\mathrm{TE}}\in\{16,\ldots,256\}$. We consider target function $f^{*}$ be a linear function, i.e., $f^{*}=\langle\beta,{\bm{x}}\rangle$ for some $\beta$ of unit norm. The top figures plots the ratio between student loss ${\mathcal{L}_{\mathrm{ST}}}$ and teacher loss $\mathcal{L}_{\mathrm{TE}}$, with varying $M_{\mathrm{TE}}$ and gradient flow training time $t$. In bottom figures, we fit student loss ${\mathcal{L}_{\mathrm{ST}}}$ as a power law function of $\mathcal{L}_{\mathrm{TE}}$. The empirical observations align with Theorem 3.1.
(a) $M_{\mathrm{ST}}=8192$, $d=16$
(b) $M_{\mathrm{ST}}=8192$, $d=32$
(c) $M_{\mathrm{ST}}=8192$, $d=64$
Figure 5: Abalation of weak-to-strong generalization in 2-layer ReLU networks. We use same setting as Figure 4 and compare the results with smaller student size $M_{\mathrm{ST}}$.
I.2 Linear ReLU Network Experiments
Figure 6 shows full experimental results for LABEL:ass:diagonalfeatcov. It shows the same conclusions as Figure 3, but with $k=10$ and $100$, providing additional verification of Theorem 3.2.
(a) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}^{2}$ v.s. $t$ at $k=1$
(b) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}^{2}$ v.s. $t$ at $k=10$
(c) ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}^{2}$ v.s. $t$ at $k=100$
Figure 6: Weak-to-Strong generalization happens in antitropic random linear feature networks (LABEL:ass:diagonalfeatcov). Here we used an input distribution as in Theorem 3.2, with $k=1,10,100$ and a target function $f^{*}=\langle\beta,{\bm{x}}\rangle$ where $\beta$ is a unit norm vector with non-zero values for first $k$ coordinates. The figures the ratio between the student loss ${\mathcal{L}_{\mathrm{ST}}}$ and squared teacher loss $\mathcal{L}_{\mathrm{TE}}^{2}$, with varying teacher size $M_{\mathrm{TE}}$, and where the dimensionality $d=M_{\mathrm{TE}}^{3/2}+k$ as set as in the scaling of Theorem 3.2, as a function of the gradient flow time $t$. With proper early stopping time ${\mathcal{L}_{\mathrm{ST}}}/\mathcal{L}_{\mathrm{TE}}^{2}$ converges to approximately $1$ as $M_{\mathrm{TE}}$ grows, confirming that for large $M_{\mathrm{TE}}$ we have ${\mathcal{L}_{\mathrm{ST}}}\propto\mathcal{L}_{\mathrm{TE}}^{2}$ as in Theorem 3.2.